### v1

# fastkqr: A Fast Algorithm for Kernel Quantile Regression

Abstract
Quantile regression is a powerful tool for robust and heterogeneous learning that has seen applications in a diverse range of applied areas. However, its broader application is often hindered by the substantial computational demands arising from the non-smooth quantile loss function. In this paper, we introduce a novel algorithm named fastkqr, which significantly advances the computation of quantile regression in reproducing kernel Hilbert spaces. The core of fastkqr is a finite smoothing algorithm that magically produces exact regression quantiles, rather than approximations. To further accelerate the algorithm, we equip fastkqr with a novel spectral technique that carefully reutilizes matrix computations. In addition, we extend fastkqr to accommodate a flexible kernel quantile regression with a data-driven crossing penalty, addressing the interpretability challenges of crossing quantile curves at multiple levels. We have implemented fastkqr in a publicly available R package. Extensive simulations and real applications show that fastkqr matches the accuracy of state-of-the-art algorithms but can operate up to an order of magnitude faster.
Keywords: finite smoothing algorithm, majorization minimization principle, non-crossing penalty, reproducing kernel Hilbert space
1 Introduction
Quantile regression (Koenker and Bassett, 1978) is a popular tool in statistics and econometrics. The method extends median regression from fitting the conditional median to modeling a suite of conditional quantile functions, providing a more comprehensive and nuanced view of the relationship between a response variable and its predictors. One of the key advantages of quantile regression, also rooted in median regression, is its robustness against outliers in the response direction. Since its introduction, quantile regression has been adapted in various research areas, including survival analysis (Peng and Huang, 2008; Wang and Wang, 2009), longitudinal data modeling (Koenker, 2004), machine learning (Meinshausen and Ridgeway, 2006; Fakoor et al., 2023), and so on, and has seen widespread applications in fields such as finance, ecology, healthcare, and engineering. For detailed introductions and the latest developments in quantile regression, see Koenker (2017) and Koenker et al. (2018).
Despite its popularity, one primary limitation of quantile regression is its high computational cost, which is also inherited from its median regression origins. This computational burden is mainly due to its non-smooth loss function. To address this, linear quantile regression is often formulated as a linear programming program and solved using the simplex method (Koenker and d’Orey, 1994) or the interior point algorithm (Portnoy and Koenker, 1997). However, the computation becomes more challenging when it comes to the kernel quantile regression (KQR, Takeuchi et al., 2006; Li et al., 2007), the method that is essential for estimating non-linear conditional quantile functions. KQR is typically solved using the interior point method, which has been implemented in the state-of-the-art R package kernlab (Karatzoglou et al., 2004), but in principle, the algorithm only provides approximate solutions to the original problem. A seminal work finding the exact solution of linear quantile regression was developed by Chen (2007). However, Chen’s algorithm works only for linear quantile regression, and extending it to KQR is not straightforward. Alternatively, one can consider the least angle regression (LARS) algorithm for computing the exact solution paths of KQR (Hastie et al., 2004; Li et al., 2007; Takeuchi et al., 2009), but it is empirically not as fast as kernlab. Recently, the kernel convolution technique has been used to smooth the quantile regression to efficiently find the approximate solutions (Fernandes et al., 2021; Tan et al., 2022; He et al., 2023).
The first main contribution of this work is the development of a fast algorithm called fastkqr to alleviate the computational burden of KQR. Our core strategy involves smoothing the original problem and recovering the exact solution by leveraging some unique properties of the quantile loss. To solve the smoothed problem, we introduce a novel spectral technique that builds upon the accelerated proximal gradient descent. With this technique, the algorithm operates at a complexity of only $\mathcal{O}(n^{2})$ after an initial eigen-decomposition of the kernel matrix. This efficient implementation makes our algorithm scalable for the KQR computation that involves numerous tuning parameters with different quantile levels.
In addition to the computational challenges mentioned above, quantile regression also poses a notable interpretability difficulty, which arises when multiple quantile functions estimated at different levels cross each other (Cole, 1988; He, 1997). This is a situation commonly encountered in practice due to finite data samples. The issue can be exemplified in a benchmark data set GAGurine from the R package MASS (Venables and Ripley, 2013). This data set records the concentration of urinary glycosaminoglycans (GAGs) for 314 children aged 0 to 17 years, with the age of the children as the predictor. As depicted in the top panel of Figure 1, five quantile curves are fitted at various levels, with crossings highlighted by gray zones where they occur. Several strategies have been proposed in the literature to address the crossing issue; examples include location-shift modeling (He, 1997), heteroscedastic location-scale modeling (Shim et al., 2009), rearranging (Chernozhukov et al., 2010), joint estimation (Sangnier et al., 2016), deep learning (Brando et al., 2022; Shen et al., 2024), and imposition of non-crossing constraints (Takeuchi et al., 2006; Bondell et al., 2010; Liu and Wu, 2011), among others.
Figure 1: Fitting quantile curves at various levels using the GAGurine data. In both panels, gray dots represent the data points. The quantile curves are fitted at levels $\tau=0.1,0.3,0.5,0.7$, and $0.9$. The top panel shows quantile curves fitted individually at each level and the gray bands highlight the zones where crossings occur. The bottom panel displays the quantile curves fitted simultaneously using the NCKQR model, where no crossings are present.
Our second major contribution addresses the crossing issue in KQR by enhancing the capabilities of fastkqr to produce non-crossing kernel quantile regression (NCKQR). Following the approach of a non-crossing constraint (Takeuchi et al., 2006; Bondell et al., 2010; Liu and Wu, 2011), we adopt a soft non-crossing penalty. This soft penalty not only eases computation but also adds flexibility, allowing the degree of crossing to be adjusted according to specific scenario tolerances. To manage the increased computational load, we fit NCKQR using a specialized majorization-minimization algorithm featuring two majorization steps. The resulting NCKQR algorithm efficiently tackles the non-crossing penalty, maintaining the same level of computational complexity as fitting a single-level KQR. To give a quick demonstration, the bottom panel of Figure 1 depicts the five quantile curves fitted using NCKQR, showcasing that no crossings occur.
Our numerical studies demonstrate that fastkqr can be significantly faster than the state-of-the-art solver, kernlab. For example, we fit KQR using fastkqr and kernlab on simulated data considered by Yuan (2006). With a sample size $n=1000$, fastkqr completed the computation in 70 seconds. In contrast, kernlab, while maintaining comparable accuracy with fastkqr, took about 700 seconds. We have implemented our algorithms in an R package fastkqr, which is publicly available on the Comprehensive R Archive Network (CRAN) at https://CRAN.R-project.org/package=fastkqr.
The rest of the paper is organized as follows. In Section 2, we review the KQR and introduce the fastkqr algorithm to solve it. In Section 3, we study the non-crossing KQR and extend fastkqr to efficiently solve the non-crossing problem. In Section 4, we present extensive numerical studies to demonstrate the competitive performance of our proposed algorithms. All technical proofs and additional experiments are provided in the online supplementary material.
2 Fast Finite Smoothing Algorithm
In this section, we propose our fastkqr algorithm. We begin with a smooth surrogate for the quantile loss function in Section 2.1 and develop the exact finite smoothing algorithm in Section 2.2. We propose an accelerated proximal gradient descent algorithm to solve the smoothed problem in Section 2.3, which is further accelerated by a fast spectral technique introduced in Section 2.4.
2.1 A Smooth surrogate for kernel quantile regression
Given $\tau\in(0,1)$, KQR estimates the $\tau$th quantile function of a response $y\in\mathbb{R}$ with covariates $\mathbf{x}\in\mathbb{R}^{p}$ through
$\min_{b\in\mathbb{R},f\in\mathcal{H}_{K}}\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b-f(\mathbf{x}_{i}))+\frac{\lambda}{2}\|{}f\|_{\mathcal{H}_{K}}^{2},$ (1)
where $\rho_{\tau}(t)=t(\tau-I(t<0))$ is the quantile check loss, $\mathcal{H}_{K}$ is the reproducing kernel Hilbert space (RKHS) induced by a kernel function $K$, and $\lambda$ is a tuning parameter governing the model complexity. A commonly used kernel function is the radial basis kernel, $K\left(\mathbf{x},\mathbf{x}^{\prime}\right)=\exp(-\frac{\|\mathbf{x}-\mathbf{x}^{\prime}\|^{2}}{2\sigma^{2}})$, where $\sigma$ is the kernel bandwidth parameter. By the representer theorem (Wahba, 1990), $\hat{f}$ has a finite-dimensional representation in terms of the kernel function, that is, $\hat{f}(\mathbf{x})=\sum_{i=1}^{n}\hat{\alpha}_{i}K\left(\mathbf{x}_{i},\mathbf{x}\right)$. Thus, the estimation of the function in problem (1) can be transformed into a finite-dimensional optimization problem,
$(\hat{b},\hat{\boldsymbol{\alpha}})=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}G(b,\boldsymbol{\alpha}),$ (2)
where $\hat{\boldsymbol{\alpha}}=(\hat{\alpha}_{1},\hat{\alpha}_{2},\ldots,\hat{\alpha}_{n})^{\top}$,
$G(b,\boldsymbol{\alpha})=\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha},$
and each $\mathbf{K}_{i}^{\top}$ is the $i$th row of $\mathbf{K}$, the $n\times{}n$ symmetric kernel matrix whose $(i,j)$th entry is $K(\mathbf{x}_{i},\mathbf{x}_{j})$.
Solving problem (2) is challenging primarily because the check loss function is non-smooth. To efficiently compute KQR, we propose to first smooth the original problem. Given $\tau\in(0,1)$, we introduce a $\gamma$-smoothed check loss function,
$H_{\gamma,\tau}(t)=\begin{cases}(\tau-1)t&\text{if}\enskip{}t<-\gamma,\\ \frac{t^{2}}{4\gamma}+t(\tau-\frac{1}{2})+\frac{\gamma}{4}&\text{if}\enskip{}-\gamma\leq t\leq\gamma,\\ \tau{}t&\text{if}\enskip{}t>\gamma,\end{cases}$ (3)
where $\gamma>0$ is the smoothing parameter. We then give the following smooth surrogate of the objective function in problem (2):
$G^{\gamma}(b,\boldsymbol{\alpha})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}.$ (4)
2.2 An exact finite smoothing algorithm
We now elaborate on how the exact KQR solution of problem (2), $(\hat{b},\hat{\boldsymbol{\alpha}})=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}G(b,\boldsymbol{\alpha})$, can be obtained based on our smoothing surrogate.
We first define $S_{0}=\{i\colon{}y_{i}=\hat{b}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}\}\subseteq\{1,2,\ldots,n\}$ to be the singular set of problem (2). Note that if the singular set $S_{0}$ is known, the following proposition shows that the exact KQR solution can be obtained from a smooth optimization problem with simple linear constraints.
Proposition 1.
Suppose $S_{0}$, the singular set of problem (2), is known. Define
$\begin{split}&(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}G^{\gamma}(b,\boldsymbol{\alpha}),\enskip{}\text{subject to }y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in{}S_{0}.\end{split}$
Then, $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=(\hat{b},\hat{\boldsymbol{\alpha}})$ holds, where $(\hat{b},\hat{\boldsymbol{\alpha}})$ is the solution of problem (2).
Although Proposition 1 suggests that the exact KQR problem can be efficiently solved through a smoothed optimization problem, it is not practically useful since the singular set $S_{0}$ is unknown before $(\hat{b},\hat{\boldsymbol{\alpha}})$ is actually obtained. To address this issue, we present a set expansion method. Specifically, for a given $\gamma>0$, we use a set $S\subseteq{}{}S_{0}$ as input and solve the following problem,
$\begin{split}&(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}G^{\gamma}(b,\boldsymbol{\alpha}),\enskip{}\text{subject to }y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in{}S.\end{split}$
Based on solution $(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})$, the set expansion method outputs a set
$\mathcal{E}(S)\equiv\{i\colon-\gamma\leq{}y_{i}-\tilde{b}^{\gamma}-\mathbf{K}_{i}^{\top}\tilde{\boldsymbol{\alpha}}^{\gamma}\leq\gamma\}.$
Define $\gamma_{0}=\min_{i\notin{}S_{0}}|y_{i}-\hat{b}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}|$, $D_{\gamma_{0}/2}=\{(b,\boldsymbol{\alpha})\colon\|{}b\mathbf{1}_{n}+\mathbf{K}\boldsymbol{\alpha}-\hat{b}\mathbf{1}_{n}-\mathbf{K}\hat{\boldsymbol{\alpha}}\|_{\infty}\geq\gamma_{0}/2\}$, $\rho=\inf_{(b,\boldsymbol{\alpha})\in{}D_{\gamma_{0}/2}}[G(b,\boldsymbol{\alpha})-G(\hat{b},\hat{\boldsymbol{\alpha}})]>0$, and $\gamma^{*}=\min\{\gamma_{0}/2,4\rho\}$. The following theorem shows that the output set $\mathcal{E}(S)$ is bounded between the input $S$ and the singular set $S_{0}$.
Theorem 2.
For any set $S\subseteq{}{}S_{0}$ and $\gamma<\gamma^{*}$, if the set expansion method outputs a set $\mathcal{E}(S)$ from $S$, then we have $S\subseteq\mathcal{E}(S)\subseteq{}S_{0}$.
Theorem 2 establishes that, by the set expansion method, any set $S\subseteq{}{}S_{0}$ will either remain unchanged or be expanded towards the singular set $S_{0}$. Therefore, we can initiate the process with the empty set $\hat{S}=\emptyset\subseteq S_{0}$ and keep expanding $\hat{S}$ by iteratively applying the set expansion method until $\hat{S}$ remains unchanged, i.e., $\hat{S}=\mathcal{E}(\hat{S})$. Since the sample size is finite, $\hat{S}$ becomes unchanged in finite steps. We thus name the whole process the finite smoothing algorithm.
The following theorem indicates that the exact KQR solution can be obtained based on $\hat{S}$. The uniqueness of the KQR solution naturally gives $\hat{S}=S_{0}$. Hence, the singular set $S_{0}$ can be constructed using the finite smoothing algorithm.
Theorem 3.
For any $\gamma\in(0,\gamma^{*})$, if there exists a set $\hat{S}\subseteq{}S_{0}$ such that $\mathcal{E}(\hat{S})=\hat{S}$, then $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=(\hat{b},\hat{\boldsymbol{\alpha}})$ holds, where
$\begin{split}&(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}G^{\gamma}(b,\boldsymbol{\alpha}),\enskip{}\text{subject to }y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in{}\hat{S},\end{split}$ (5)
and $(\hat{b},\hat{\boldsymbol{\alpha}})$ is the solution to problem (2).
Therefore, with the finite smoothing algorithm, the exact solution of problem (2) can be obtained by iteratively solving problem (5) and augmenting the set $\hat{S}$. Since $\gamma^{*}$ is still unknown in practice, we handle this by repeatedly implementing the above procedure with a decreasing sequence of values of $\gamma$. The algorithm is terminated once a solution satisfies the Karush–Kuhn–Tucker (KKT) conditions of problem (2). In our implementation, we initiate this process with $\gamma=1$ and iteratively update it by reducing $\gamma$ to a quarter of its previous value, i.e., $\gamma\leftarrow\gamma/4$. This approach generally converges after only three or four iterations of updating $\gamma$.
2.3 Solving the smoothed kernel quantile regression
In this section, we develop an accelerated proximal gradient descent (APGD) algorithm to solve problem (5). We first consider the unconstrained optimization, say, problem (5) with $\hat{S}=\emptyset$.
Let $(b^{(1)},\boldsymbol{\alpha}^{(1)})$ be the initial value of $(b,\boldsymbol{\alpha})$. For each $k=1,2,\ldots$, let $\mathbf{z}^{(k)}$ be the $n$-vector with the $i$th element $H^{\prime}_{\gamma,\tau}(y_{i}-b^{(k)}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)})$ and compute
$\left(\begin{array}[]{c}b^{(k+1)}\\ \boldsymbol{\alpha}^{(k+1)}\end{array}\right)$ $=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\frac{1}{4n\gamma}\left\|b\mathbf{1}+\mathbf{K}\boldsymbol{\alpha}-(b^{(k)}\mathbf{1}{}+\mathbf{K}\boldsymbol{\alpha}^{(k)}+{}2\gamma\mathbf{z}^{(k)})\right\|_{2}^{2}+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}$ (6)
$=\left(\begin{array}[]{c}b^{(k)}\\ \boldsymbol{\alpha}^{(k)}\end{array}\right)+2\gamma\mathbf{P}_{\gamma,\lambda}^{-1}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}-n\lambda\mathbf{K}\boldsymbol{\alpha}^{(k)}\end{array}\right),$
where
$\mathbf{P}_{\gamma,\lambda}=\left(\begin{array}[]{cc}n&\mathbf{1}^{\top}\mathbf{K}\\ \mathbf{K}^{\top}\mathbf{1}&\mathbf{K}^{\top}\mathbf{K}+2n\gamma\lambda\mathbf{K}\end{array}\right).$
The step size is directly obtained from the following Lipschitz condition of $H_{\gamma,\tau}^{\prime}(t)$,
$|H_{\gamma,\tau}^{\prime}(c_{1})-H_{\gamma,\tau}^{\prime}(c_{2})|\leq\frac{1}{2\gamma}|c_{1}-c_{2}|,\ \forall c_{1},c_{2}\in\mathbb{R}.$
In fact, the update in equation (6) is essentially equivalent to performing a step of the majorization-minimization (MM) algorithm (Hunter and Lange, 2000),
$\begin{split}\left(\begin{array}[]{c}b^{(k+1)}\\ \boldsymbol{\alpha}^{(k+1)}\end{array}\right)=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}&\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b^{(k)}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)})\\ &+\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}^{\prime}(y_{i}-b^{(k)}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)})(b^{(k)}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})\\ &+\frac{1}{4n\gamma}\lVert{}b^{(k)}\mathbf{1}+\mathbf{K}\boldsymbol{\alpha}^{(k)}-b\mathbf{1}-\mathbf{K}\boldsymbol{\alpha}\rVert_{2}^{2}+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha},\end{split}$
whose objective function majorizes the function $G^{\gamma}(b,\boldsymbol{\alpha})$.
We further apply Nesterov’s acceleration (Nesterov, 1983; Beck and Teboulle, 2009) to boost the algorithm. Given a sequence $\{c_{k}\}_{k\geq 1}$, such that $c_{1}=1$ and $c_{k+1}=1/2+(1+4c_{k}^{2})^{1/2}/2$ for $k\geq 1$, let $(b^{(0)},{\boldsymbol{\alpha}^{(0)}})$ and $(b^{(1)},{\boldsymbol{\alpha}^{(1)}})$ be the first two iterates. For each $k=1,2,\ldots$, we solve $(b^{(k+1)},\boldsymbol{\alpha}^{(k+1)})$ from the following problem
$\begin{split}\left(\begin{array}[]{c}b^{(k+1)}\\ \boldsymbol{\alpha}^{(k+1)}\end{array}\right)&=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\frac{1}{4n\gamma}\left\|b\mathbf{1}+\mathbf{K}\boldsymbol{\alpha}-(\bar{b}^{(k)}\mathbf{1}+\mathbf{K}\boldsymbol{\bar{\alpha}}^{(k)}+2\gamma\bar{\mathbf{z}}^{(k)})\right\|_{2}^{2}+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}\\ &=\left(\begin{array}[]{c}\bar{b}^{(k)}\\ \bar{\boldsymbol{\alpha}}^{(k)}\end{array}\right)+2\gamma\mathbf{P}_{\gamma,\lambda}^{-1}\bar{\boldsymbol{\zeta}},\end{split}$ (7)
where
$\left(\begin{array}[]{c}\bar{b}^{(k)}\\ \boldsymbol{\bar{\alpha}}^{(k)}\end{array}\right)=\left(\begin{array}[]{c}b^{(k)}\\ \boldsymbol{\alpha}^{(k)}\end{array}\right)+\left(\frac{c_{k}-1}{c_{k+1}}\right)\left(\begin{array}[]{c}b^{(k)}-b^{(k-1)}\\ \boldsymbol{\alpha}^{(k)}-\boldsymbol{\alpha}^{(k-1)}\end{array}\right)$
and
$\bar{\boldsymbol{\zeta}}=\left(\begin{array}[]{c}\mathbf{1}^{\top}\bar{\mathbf{z}}^{(k)}\\ \mathbf{K}^{\top}\bar{\mathbf{z}}^{(k)}-n\lambda\mathbf{K}\bar{\boldsymbol{\alpha}}^{(k)}\end{array}\right)\enskip\text{with}\enskip{}\bar{z}_{i}^{(k)}=H_{\gamma,\tau}^{\prime}(y_{i}-\bar{b}^{(k)}-\mathbf{K}_{i}^{\top}\boldsymbol{\bar{\alpha}}^{(k)}),\,i=1,\ldots,n.$
The standard theory of the APGD algorithm gives the following convergence analysis.
Proposition 4.
Suppose $G^{\gamma}(b,\alpha)$ is defined in equation (4), $(b^{*},\boldsymbol{\alpha}^{*})$ is the global minimizer, and $(b^{(k)},\boldsymbol{\alpha}^{(k)})$ is the solution at the $k$th iteration of the APGD algorithm. It holds that
$G^{\gamma}(b^{(k)},\boldsymbol{\alpha}^{(k)})-G^{\gamma}(b^{*},\boldsymbol{\alpha}^{*})\leq\dfrac{1}{\gamma k^{2}}\left((b^{(0)}-b^{*})^{2}+\|\mathbf{K}(\boldsymbol{\alpha}^{(0)}-\boldsymbol{\alpha}^{*})\|_{2}^{2}\right).$
We now address the constraint in problem (5), for which we consider the projected gradient descent algorithm. Specifically, after obtaining $(b^{(k)},\boldsymbol{\alpha}^{(k)})$ for each $k$, we project the solution onto the feasible region associated with the constraint by solving the following optimization problem,
$(\tilde{b},\tilde{\boldsymbol{\alpha}})=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}(b-b^{(k)})^{2}+\|\mathbf{K}(\boldsymbol{\alpha}-\boldsymbol{\alpha}^{(k)})\|_{2}^{2},\enskip{}\text{subject to }y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},i\in\hat{S}.$ (8)
It can be shown that the solution is $\tilde{b}=b^{(k)}+\frac{1}{|\hat{S}|+1}\sum_{i\in\hat{S}}[y_{i}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)}]$ and $\tilde{\boldsymbol{\alpha}}=\mathbf{K}^{-1}\boldsymbol{\theta}$, where $|\hat{S}|$ denotes the number of elements in the set $\hat{S}$, and $\boldsymbol{\theta}\in\mathbb{R}^{n}$ with $\theta_{i}=y_{i}-\tilde{b}$ if $i\in\hat{S}$ and $\theta_{i}=\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)}$ otherwise. Subsequently, we use $(\tilde{b},\tilde{\boldsymbol{\alpha}})$ in place of $(b^{(k)},\boldsymbol{\alpha}^{(k)})$ in problem (6) to move the APGD algorithm forward to obtain $(b^{(k+1)},\boldsymbol{\alpha}^{(k+1)})$. In practice, we find that the performance is often nearly identical if the above projection is applied only once to the unconstrained solution of problem (5). This implementation is effective mainly because the solution of the unconstrained problem barely violates the constraint when $\gamma$ is sufficiently small.
2.4 A fast spectral technique
We note that the computational bottleneck of the APGD algorithm discussed in Section 2.3 is the inversion of $\mathbf{P}_{\gamma,\lambda}$, which typically has a computational complexity of $\mathcal{O}(n^{3})$. The whole KQR algorithm can become very expensive because the matrix inversion must be repeated for every $\mathbf{P}_{\gamma,\lambda}$ as the smoothing parameter $\gamma$ and the tuning parameter $\lambda$ vary. Although we could consider alternative algorithms, such as gradient descent or quasi-Newton methods, to circumvent matrix inversion, our empirical studies indicate that the precision of these alternatives is generally inferior to that of the APGD algorithm, unless they are executed for an excessive number of iterations.
To accelerate the APGD algorithm, we develop a spectral technique. We begin with the eigen-decomposition of the kernel matrix, $\mathbf{K}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$, where $\boldsymbol{\Lambda}$ is diagonal and $\mathbf{U}$ is orthogonal. Note that this step is free of the parameters $\gamma$ and $\lambda$. Define $\boldsymbol{\Pi}_{\gamma,\lambda}=\boldsymbol{\Lambda}^{2}+2n\gamma\lambda\boldsymbol{\Lambda}$, $\mathbf{v}=\mathbf{U}\boldsymbol{\Lambda}\Pi_{\gamma,\lambda}^{-1}\mathbf{U}^{\top}\mathbf{1}$, and $g=1/\bigl{(}n\mathbf{1}^{\top}\mathbf{U}\boldsymbol{\Lambda}\Pi_{\gamma,\lambda}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\mathbf{1}\bigr{)}$. We have
$\begin{split}\mathbf{P}_{\gamma,\lambda}^{-1}&=\left(\begin{array}[]{cc}n&\mathbf{1}^{\top}\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}\\ \mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}\mathbf{1}&\mathbf{U}\Pi_{\gamma,\lambda}\mathbf{U}^{\top}\end{array}\right)^{-1}=g\Biggl{(}\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\Biggr{)}\Bigl{(}\begin{array}[]{ll}1&-\mathbf{v}^{\top}\end{array}\Bigr{)}+\left(\begin{array}[]{cc}0&\mathbf{0}^{\top}\\ \mathbf{0}&\mathbf{U}\Pi_{\lambda}^{-1}\mathbf{U}^{\top}\end{array}\right).\end{split}$ (9)
Although it may initially appear that the computation of $\mathbf{P}_{\gamma,\lambda}^{-1}$ using equation (9) still has the complexity of $\mathcal{O}(n^{3})$, as matrix multiplications are still involved, it is important to note that the APGD update requires only the direct computation of $\mathbf{P}_{\gamma,\lambda}^{-1}\boldsymbol{\zeta}$ rather than the matrix inversion itself. Thus equation (9) gives
$\mathbf{P}_{\gamma,\lambda}^{-1}\boldsymbol{\zeta}=g\left\{\mathbf{1}^{\top}\mathbf{z}-\mathbf{v}^{\top}\mathbf{K}\left(\mathbf{z}+n\lambda\boldsymbol{\alpha}\right)\right\}\Biggl{(}\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\Biggr{)}+\left(\begin{array}[]{c}0\\ \mathbf{U}\boldsymbol{\Pi}_{\gamma,\lambda}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\left(\mathbf{z}+n\lambda\boldsymbol{\alpha}\right)\end{array}\right),$ (10)
of which the computational complexity is only $\mathcal{O}\left(n^{2}\right)$ when computed from right to left, that is, by only performing matrix-vector multiplications.
In our implementation, we use a warm-start strategy to further amplify the effect of the spectral technique. Specifically, we solve problem (2) with a sequence of tuning parameters $\lambda^{[1]},\lambda^{[2]},\ldots,\lambda^{[L]}$ to obtain the corresponding solutions $(\hat{b}^{[1]},\hat{\boldsymbol{\alpha}}^{[1]}),(\hat{b}^{[2]},\hat{\boldsymbol{\alpha}}^{[2]}),\ldots,(\hat{b}^{[L]},\hat{\boldsymbol{\alpha}}^{[L]})$. With $l>1$, $(\hat{b}^{[l-1]},\hat{\boldsymbol{\alpha}}^{[l-1]})$ is summoned to initialize the finite smoothing and APGD algorithms to solve $(\hat{b}^{[l]},\hat{\boldsymbol{\alpha}}^{[l]})$. Therefore, thanks to both the warm start and the spectral technique, except for the only step of eigen-decomposition, costing the $\mathcal{O}(n^{3})$ complexity, the rest of the entire KQR algorithm to solve problem (2) is only $\mathcal{O}(n^{2})$, hence the speed is significantly enhanced.
The algorithm fastkqr is summarized in Algorithm 1 in the online supplemental material.
3 Non-crossing Kernel Quantile Regression
In this section, we propose a non-crossing kernel quantile regression (NCKQR) method to address the crossing issue of the quantile curves fitted at various levels.
3.1 Methodology
When KQR is fitted at multiple quantile levels individually, say, for $10<\tau_{1}<\tau_{2}<\cdots<\tau_{T}<$ 1,
$\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau_{t}}\left(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}\right)+\frac{\lambda}{2}\boldsymbol{\alpha}_{\tau_{t}}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}},\ t=1,2,\ldots,T,$
the fitted curves may cross each other. To avoid the occurrence of crossing, Takeuchi et al. (2006), Bondell et al. (2010), and Liu and Wu (2011) consider fitting all the quantile curves simultaneously with a hard non-crossing constraint,
$\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}}$ $\sum_{t=1}^{T}\left[\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau_{t}}\left(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}\right)+\frac{\lambda}{2}\boldsymbol{\alpha}_{\tau_{t}}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]$ (11)
subject to $b_{\tau_{t_{1}}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t_{1}}}<b_{\tau_{t_{2}}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t_{2}}}\text{ for all }t_{1}<t_{2}\text{ and }i=1,2,\ldots,n,$
where the notation $\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}$ represents the collection of $b_{\tau_{1}},\boldsymbol{\alpha}_{\tau_{1}},b_{\tau_{2}},\boldsymbol{\alpha}_{\tau_{2}},\cdots,b_{\tau_{T}},\boldsymbol{\alpha}_{\tau_{T}}$ for ease of presentation. In the sequel, we extend this notation to $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$ and $\{\tilde{b}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$ to represent their respective counterparts.
With the hard constraint imposed in problem (11), the quantile regression curves fitted on finite-sample data do not cross. However, the inequality constraints introduced by this formulation may largely increase the computational cost. As such, we propose using a soft crossing penalty. This approach does not have any inequality constraint. Moreover, it provides practitioners with some flexibility to tolerate a certain level of crossing. Specifically, our NCKQR is defined as:
$\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}=\operatorname*{argmin}_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}}Q\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\right),$ (12)
where
$Q\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\right)=$ $\sum_{t=1}^{T}\left[\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau_{t}}\left(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}\right)\right]+\sum_{t=1}^{T}\frac{\lambda_{2}}{2}\boldsymbol{\alpha}^{\top}_{\tau_{t}}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}$
$+\lambda_{1}\sum_{t=1}^{T-1}\left[\sum_{i=1}^{n}V\left(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}}\right)\right],$
and $V$ is the soft crossing penalty that is non-decreasing and $V(0)=0$. When $\lambda_{1}=0$, NCKQR reduces to fitting KQR individually at different levels; when $\lambda_{1}\to\infty$, the soft crossing penalty functions as the hard constraint to restrict the solution to the feasible set of problem (11).
A potential choice of $V$ is a rectified linear unit (ReLU), i.e., $V(t)=\max\{t,0\}$, to enforce the regularization on crossing. However, the ReLU penalty is non-smooth, which makes the computation unnecessarily challenging. Although it seems to be a possible solution that employing the smoothing algorithm developed in Section 2.1 to simultaneously smooth both the functions $\rho_{\tau_{t}}$ and $V$, additional efforts are expected to ensure the exact solution is obtained.
Yet, we may alleviate the computational burden brought by the ReLU penalty from a different angle. The use of $\rho_{\tau}$ and the pursuit of an exact solution are imperative due to the fundamental nature of quantile regression; deviating from this would essentially lead to a different statistical problem. In contrast, the use of a non-smooth ReLU penalty is not obligatory, as its smoothed counterpart can function alone as a soft crossing penalty. Thus, we opt to define NCKQR directly by employing a smooth ReLU penalty in Problem (12), formulated as follows:
$V(t)=\begin{cases}0&\text{if}\enskip{}t<-\eta,\\ \frac{t^{2}}{4\eta}+\frac{t}{2}+\frac{\eta}{4}&\text{if}\enskip{}-\eta\leq t\leq\eta,\\ t&\text{if}\enskip{}t>\eta,\end{cases}$
where $\eta=10^{-5}$ is used in our implementation. In the next section, we shall focus on finding an algorithm to obtain the exact solution of problem (12) in which $V$ is the smooth ReLU penalty.
Note that the primary focus of this paper is on the fast computation of KQR. Although the soft crossing penalty could be advocated for its additional flexibility, a theoretical investigation of this penalty is beyond the scope of this paper. Our main rationale for proposing the soft non-crossing penalty is due to computational considerations. Practitioners may directly use a large $\lambda_{1}$ value to approximate the results that would be obtained with a hard non-crossing constraint.
3.2 Exact finite smoothing algorithm for NCKQR
We now extend fastkqr to solve NCKQR. With the check loss replaced by the $\gamma$-smoothed check loss $H_{\gamma,\tau}$, we formulate a smooth surrogate objective function,
$Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\right)=$ $\sum_{t=1}^{T}\left[\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau_{t}}\left(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}\right)\right]+\sum_{t=1}^{T}\frac{\lambda_{2}}{2}\boldsymbol{\alpha}^{\top}_{\tau_{t}}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}$ (13)
$+\lambda_{1}\sum_{t=1}^{T-1}\left[\sum_{i=1}^{n}V\left(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}}\right)\right].$
According to $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$, the solution of problem (12), we first construct the singular sets, $S_{0,t}=\{i:y_{i}=\hat{b}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}\subseteq\{1,\ldots,n\}$, for each quantile level $\tau_{1},\tau_{2},\ldots,\tau_{T}$. In the following proposition, we demonstrate that $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$ can be obtained by solving the above smooth optimization problem with linear constraints associated with the singular sets. When $T=1$, the result reduces to Proposition 1 for a single-level KQR problem.
Proposition 5.
Suppose $S_{0,t}$ is known for each $t=1,2,\ldots,T$. Define
$\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}_{t=1}^{T}=\operatorname*{argmin}_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}}Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\right),\enskip{}\text{subject to}\enskip{}y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,\forall i\in{}S_{0,t},1\leq t\leq T.$
It holds that $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}_{t=1}^{T}=\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$, where $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$ is the solution to problem (12).
Similar to Proposition 1 for the single-level KQR, Proposition 5 cannot be applied in practice, since none of the $S_{0,t}$’s are known. We hereby develop a multiple-level set expansion method. For a given $\gamma>0$, we use a collection of sets $S_{1},S_{2},\ldots,S_{T}$ as inputs and proceed to solve the following optimization problem,
$\{\tilde{b}_{\tau_{t}}^{\gamma},\tilde{\boldsymbol{\alpha}}_{\tau_{t}}^{\gamma}\}_{t=1}^{T}=\operatorname*{argmin}_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}}Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\right),\enskip{}\text{subject to}\enskip{}y_{i}=\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}+b_{\tau_{t}},\,i\in{}S_{t},1\leq t\leq T.$
Then the multiple-level set expansion method outputs a collection of the following $T$ sets,
$\mathcal{E}_{t}(S_{1},S_{2},\ldots,S_{T})\equiv\{i:-\gamma{}\leq{}y_{i}-\tilde{b}_{\tau_{t}}^{\gamma}-\mathbf{K}_{i}^{\top}\tilde{\boldsymbol{\alpha}}_{\tau_{t}}^{\gamma}\leq\gamma\},\ t=1,2,\ldots,T.$
To bound these output sets, some quantities are required. For each quantile level $\tau_{t}$, we define
$\gamma_{0,t}$ $=\min_{i\notin{}S_{0,t}}|y_{i}-\hat{b}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}_{\tau_{t}}|,$
$D_{\gamma_{0},t}$ =$=\{(b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}})\colon\|b_{\tau_{t}}\mathbf{1}_{n}+\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}-\hat{b}_{\tau_{t}}\mathbf{1}_{n}-\mathbf{K}\hat{\boldsymbol{\alpha}}_{\tau_{t}}\|_{\infty}\geq\gamma_{0,t}/2\}.$
Denote by $\mathcal{D}$ the collection of $\{(b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}})\}_{t=1}^{T}$ such that $(b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}})\in D_{\gamma_{0},t}$ for each $t=1,2,\ldots,T$. Define $\gamma^{*}=\min\{\frac{1}{2}\gamma_{0,1},\frac{1}{2}\gamma_{0,2},\ldots,\frac{1}{2}\gamma_{0,T},\frac{4}{T}\rho\}$, where
$\rho=\inf_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\in\mathcal{D}}Q\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\right)-Q\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}\right)>0.$
The following theorem bounds the output sets from the multiple-level set expansion method.
Theorem 6.
With some $\gamma<\gamma^{*}$, if $S_{t}\subseteq S_{0,t}$ holds for every $t=1,2,\ldots,T$, then the output sets from the multiple-level set expansion method must satisfy
$S_{t}\subseteq\mathcal{E}_{t}(S_{1},S_{2},\ldots,S_{T})\subseteq S_{0,t},\ \forall t=1,2,\ldots,T.$
Knowing both the lower and upper bounds of the output sets, we have the following theorem showing the exact solution of problem (12) can be obtained by the finite smoothing algorithm.
Theorem 7.
For any $\gamma\in(0,\gamma^{*})$, if there exist $T$ sets, $\hat{S}_{1},\hat{S}_{2},\ldots,\hat{S}_{T}$, satisfying
  1. (1)
    $\hat{S}_{t}\in S_{0,t},\ \forall t$,
  2. (2)
    $\mathcal{E}_{t}(\hat{S}_{1},\hat{S}_{2},\ldots,\hat{S}_{T})=\hat{S}_{t},\ \forall t$,
then the solution to problem (12) can be obtained from the following optimization problem,
$\{\hat{b}_{\tau_{t}}^{\gamma},\hat{\boldsymbol{\alpha}}_{\tau_{t}}^{\gamma}\}_{t=1}^{T}=\operatorname*{argmin}_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}}Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\right),\enskip{}\text{subject to}\enskip{}y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,\forall i\in{}\hat{S}_{t},1\leq t\leq T.$ (14)
In other words, we have $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}=\{\hat{b}_{\tau_{t}}^{\gamma},\hat{\boldsymbol{\alpha}}_{\tau_{t}}^{\gamma}\}_{t=1}^{T}$.
Therefore, similar to the algorithm for the single-level KQR, Theorem 7 enables one to apply the finite smoothing algorithm to derive the exact NCKQR solution of problem (12). Starting with $\hat{S}_{t}=\emptyset\subseteq S_{0,t}$, for each $t=1,2,\ldots,T$, the multiple-level set expansion method is iteratively applied on these $T$ sets until no further changes occur. According to Theorem 7, the NCKQR solution is then obtained, provided that $\gamma<\gamma^{*}$. As $\gamma^{*}$ is unknown, to ensure that $\gamma$ is adequately small, the finite smoothing algorithm is repeated with a decreasing sequence of $\gamma$ values until a solution satisfying the KKT conditions of problem (12) is eventually identified.
3.3 Computation
In the previous section, we extended the finite smoothing algorithm to address the NCKQR problem. However, even when NCKQR is smoothed into problem (14), the non-crossing penalty increases the computational demands of NCKQR compared to those of the single-level KQR. There are three main reasons. First, the smooth ReLU and $\gamma$-smoothed quantile loss functions in problem (14) have different Lipschitz constants, hence it is challenging to determine the step size for the optimization algorithm. Second, the different Lipschitz constants in the two functions alter the update formula of the proximal gradient descent algorithm, making the direct implementation of the fast spectral technique introduced in Section 2.4 impractical. Third, even if the fast spectral technique could be utilized, the counterpart of the matrix $\mathbf{P}_{\gamma,\lambda}$ in equation (10) would be $L$ times larger in NCKQR, making the matrix operations exceedingly costly.
We first solve problem (14) without the linear constraints, say, $\hat{S}_{t}=\emptyset$, for $t=1,2,\ldots,T$. To circumvent the three challenges discussed above, we propose a specialized MM algorithm with two majorization steps. This MM algorithm can calibrate the Lipschitz constants, making the fast spectral technique feasible; it can also transform the counterpart of $\mathbf{P}_{\gamma,\lambda}$ in NCKQR into a block diagonal matrix, which effectively makes the computation scalable.
The first majorization manages the different Lipschitz constants in $V$ and $H_{\gamma,\tau}$.
Recall that $H_{\gamma,\tau}^{\prime}$ is Lipschitz continuous with constant $\gamma^{-1}$, so for any $c_{1}\neq c_{2}$, we have
$H_{\gamma,\tau}(c_{1})\leq{}H_{\gamma,\tau}(c_{2})+H_{\gamma,\tau}^{\prime}(c_{2})(c_{1}-c_{2})+\frac{1}{2\gamma}(c_{1}-c_{2})^{2}.$ (15)
Also by the definition of $V$, we can see that $V^{\prime}$ is Lipschitz continuous with constant $\eta^{-1}$. To calibrate the two different Lipschitz constants, we require $\gamma\leq\eta$, which gives
$V(c_{1})$ $\leq{}V(c_{2})+V^{\prime}(c_{2})(c_{1}-c_{2})+\frac{1}{2\eta}(c_{1}-c_{2})^{2}$ (16)
$\leq V(c_{2})+V^{\prime}(c_{2})(c_{1}-c_{2})+\frac{1}{2\gamma}(c_{1}-c_{2})^{2}.$
Let $\{b_{\tau_{t}}^{(1)},\boldsymbol{\alpha}_{\tau_{t}}^{(1)}\}_{t=1}^{T}$ be the initial value. For each $k=1,2,\ldots$ and each $t=1,2,\ldots,T$, let
$\boldsymbol{\zeta}_{\tau_{t},\lambda_{1}}^{(k)}$ $=\left(\begin{array}[]{c}\lambda_{1}\mathbf{1}^{\top}(\mathbf{q}_{\tau_{t-1}}^{(k)}-\mathbf{q}_{\tau_{t}}^{(k)})\\ \lambda_{1}\mathbf{K}^{\top}(\mathbf{q}_{\tau_{t-1}}^{(k)}-\mathbf{q}_{\tau_{t}}^{(k)})\end{array}\right)\enskip{\text{and}}\enskip\boldsymbol{\zeta}_{\tau_{t},\lambda_{2}}^{(k)}=\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}_{\tau_{t}}^{(k)}\\ \mathbf{K}^{\top}\mathbf{z}_{\tau_{t}}^{(k)}-n\lambda_{2}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}^{(k)}\end{array}\right),$
where $\mathbf{z}_{\tau_{t}}^{(k)}$ is an $n$-vector whose $i$th element is $H^{\prime}_{\gamma,\tau}(y_{i}-b_{\tau_{t}}^{(k)}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}^{(k)})$, $\mathbf{q}_{\tau_{t}}^{(k)}$ is an $n$-vector whose $i$th element is $V^{\prime}(b_{\tau_{t}}^{(k)}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}^{(k)}-b_{\tau_{t+1}}^{(k)}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}}^{(k)})$, and $\mathbf{q}_{\tau_{t}}^{(k)}=\mathbf{0}$ if $t=0$ or $t>T$. Define
$R\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\right)=$ $\sum_{t=1}^{T}\left[\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau_{t}}\left(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}\right)\right]$
$+\lambda_{1}\sum_{t=1}^{T-1}\left[\sum_{i=1}^{n}V\left(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}}\right)\right].$
With $Q^{\gamma}$ defined in equation (13), according to inequalities (15) and (16), we have
$Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\right)=\sum_{t=1}^{T}\frac{\lambda_{2}}{2}\boldsymbol{\alpha}^{\top}_{\tau_{t}}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}+R\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\right)$ (17)
$\leq$ $\sum_{t=1}^{T}\frac{\lambda_{2}}{2}\boldsymbol{\alpha}^{\top}_{\tau_{t}}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}+R\left(\{b_{\tau_{t}}^{(k)},\boldsymbol{\alpha}_{\tau_{t}}^{(k)}\}_{t=1}^{T}\right)$
$+\sum_{t=1}^{T}\left[-\boldsymbol{\zeta}_{\tau_{t},\lambda_{2}}^{\top}\left(\begin{array}[]{c}b_{\tau_{t}}-b^{(k)}_{\tau_{t}}\\ \boldsymbol{\alpha}_{\tau_{t}}-\boldsymbol{\alpha}^{(k)}_{\tau_{t}}\end{array}\right)+\frac{1}{4\gamma}\left(\begin{array}[]{c}b_{\tau_{t}}-b^{(k)}_{\tau_{t}}\\ \boldsymbol{\alpha}_{\tau_{t}}-\boldsymbol{\alpha}^{(k)}_{\tau_{t}}\end{array}\right)^{\top}\mathbf{P}_{\gamma,\lambda_{2}}\left(\begin{array}[]{c}b_{\tau_{t}}-b^{(k)}_{\tau_{t}}\\ \boldsymbol{\alpha}_{\tau_{t}}-\boldsymbol{\alpha}^{(k)}_{\tau_{t}}\end{array}\right)\right]$
$+\sum_{t=1}^{T}\left[\boldsymbol{\zeta}_{\tau_{t},\lambda_{1}}^{\top}\left(\begin{array}[]{c}b_{\tau_{t}}-b^{(k)}_{\tau_{t}}\\ \boldsymbol{\alpha}_{\tau_{t}}-\boldsymbol{\alpha}^{(k)}_{\tau_{t}}\end{array}\right)\right]+\frac{1}{4\gamma}\left(\begin{array}[]{c}b_{\tau_{1}}-b^{(k)}_{\tau_{1}}\\ \boldsymbol{\alpha}_{\tau_{1}}-{\boldsymbol{\alpha}}^{(k)}_{\tau_{1}}\\ \ldots\\ b_{\tau_{T}}-b^{(k)}_{\tau_{T}}\\ \boldsymbol{\alpha}_{\tau_{T}}-{\boldsymbol{\alpha}}^{(k)}_{\tau_{T}}\end{array}\right)^{\top}{\boldsymbol{\Phi}}_{\gamma,\lambda_{1},\lambda_{2}}\left(\begin{array}[]{c}b_{\tau_{1}}-b^{(k)}_{\tau_{1}}\\ \boldsymbol{\alpha}_{\tau_{1}}-{\boldsymbol{\alpha}}^{(k)}_{\tau_{1}}\\ \ldots\\ b_{\tau_{T}}-b^{(k)}_{\tau_{T}}\\ \boldsymbol{\alpha}_{\tau_{T}}-{\boldsymbol{\alpha}}^{(k)}_{\tau_{T}}\\ \end{array}\right)$
$\equiv$ $\tilde{Q}_{\mathrm{M}}^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}\right),$
where
$\mathbf{P}_{\gamma,\lambda_{2}}=\left(\begin{array}[]{cc}n&\mathbf{1}^{\top}\mathbf{K}\\ \mathbf{K}^{\top}\mathbf{1}&\mathbf{K}^{\top}\mathbf{K}+2n\gamma\lambda_{2}\mathbf{K}\end{array}\right)$
and a Block Toeplitz matrix ${\boldsymbol{\Phi}}_{\gamma,\lambda_{1},\lambda_{2}}\in\mathbb{R}^{(n+1)(T-1)\times(n+1)(T-1)}$ such that
${\boldsymbol{\Phi}}_{\gamma,\lambda_{1},\lambda_{2}}=\left(\begin{array}[]{ccccc}\mathbf{B}&-\mathbf{B}&&\\ -\mathbf{B}&2\mathbf{B}&-\mathbf{B}&\\ &\ddots&\ddots&\ddots\\ &&-\mathbf{B}&2\mathbf{B}&-\mathbf{B}\\ &&&-\mathbf{B}&\mathbf{B}\end{array}\right)\enskip\text{with}\enskip\mathbf{B}=\left(\begin{array}[]{cc}n\lambda_{1}&\lambda_{1}\mathbf{1}^{\top}\mathbf{K}\\ \lambda_{1}\mathbf{K}^{\top}\mathbf{1}&\lambda_{1}\mathbf{K}^{\top}\mathbf{K}\end{array}\right).$
In order to efficiently solve $\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T}$ using inequality (17), we note the matrix operations directly involving ${\boldsymbol{\Phi}}_{\gamma,\lambda_{1},\lambda_{2}}$ can be computationally prohibitive. We thus propose to employ a second majorization to craft a block diagonal ${\boldsymbol{\Psi}}_{\gamma,\lambda_{1},\lambda_{2}}$ that majorizes ${\boldsymbol{\Phi}}_{\gamma,\lambda_{1},\lambda_{2}}$,
${\boldsymbol{\Psi}}_{\gamma,\lambda_{1},\lambda_{2}}={\boldsymbol{\Phi}}_{\gamma,\lambda_{1},\lambda_{2}}+\left(\begin{array}[]{ccccc}\mathbf{C}&\mathbf{B}&&\\ \mathbf{B}&\mathbf{C}&\mathbf{B}&\\ &\ddots&\ddots&\ddots&\\ &&\mathbf{B}&\mathbf{C}&\mathbf{B}\\ &&&\mathbf{B}&\mathbf{C}\end{array}\right)=\left(\begin{array}[]{ccccc}\mathbf{B+C}&&&\\ &\mathbf{2B+C}&&\\ &&\ddots&&\\ &&&\mathbf{2B+C}&\\ &&&&\mathbf{B+C}\end{array}\right),$
where, with $\varepsilon$ set to $10^{-3}$,
$\mathbf{C}=\left(\begin{array}[]{cc}2n\lambda_{1}+\varepsilon\lambda_{1}&2\lambda_{1}\mathbf{1}^{\top}\mathbf{K}\\ 2\lambda_{1}\mathbf{K}^{\top}\mathbf{1}&2\lambda_{1}\mathbf{K}^{\top}\mathbf{K}+\varepsilon\mathbf{I}\end{array}\right).$
We define ${Q}^{\gamma}_{\mathrm{M}}(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T})$ by replacing the term ${\boldsymbol{\Phi}}_{\gamma,\lambda_{1},\lambda_{2}}$ in $\tilde{Q}^{\gamma}_{\mathrm{M}}(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T})$ with ${\boldsymbol{\Psi}}_{\gamma,\lambda_{1},\lambda_{2}}$. We then derive the MM algorithm to obtain $\{b_{\tau_{t}}^{(k+1)},\boldsymbol{\alpha}_{\tau_{t}}^{(k+1)}\}_{t=1}^{T}$ by minimizing ${Q}_{\mathrm{M}}^{\gamma}(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T})$. Setting the gradients of ${Q_{\mathrm{M}}^{\gamma}}(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}_{t=1}^{T})$ to be $\mathbf{0}$, we have
$\left(\begin{array}[]{c}b^{(k+1)}_{\tau_{t}}\\ \boldsymbol{\alpha}^{(k+1)}_{\tau_{t}}\end{array}\right)=\left(\begin{array}[]{c}b^{(k)}_{\tau_{t}}\\ \boldsymbol{\alpha}^{(k)}_{\tau_{t}}\end{array}\right)+2\gamma\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\boldsymbol{\varrho}^{(k)},$ (18)
where
$\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}=\left(\begin{array}[]{cc}n+4\lambda_{1}n^{2}+\varepsilon\lambda_{1}n&(4\lambda_{1}n+1)\mathbf{1}^{\top}\mathbf{K}\\ (4\lambda_{1}n+1)\mathbf{K}^{\top}\mathbf{1}&(4\lambda_{1}n+1)\mathbf{K}^{\top}\mathbf{K}+2n\gamma\lambda_{2}\mathbf{K}+\lambda_{1}\varepsilon n\end{array}\right)$
and
$\boldsymbol{\varrho}^{(k)}=\left\{\begin{array}[]{ll}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{K}\mathbf{q}^{(k)}_{\tau_{t}}\end{array}\right),&t=1,\\ \left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{K}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{K}\mathbf{q}^{(k)}_{\tau_{t-1}}\end{array}\right),&2\leq t\leq T-1,\\ \left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{K}\mathbf{q}^{(k)}_{\tau_{t-1}}\end{array}\right),&t=T.\end{array}\right.$
The fast spectral technique can be extended to address the repeated computation of inverting $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}\in\mathbb{R}^{(n+1)\times(n+1)}$, for varying values of $\gamma$, $\lambda_{1}$ and $\lambda_{2}$. Similar to the single-level KQR, we begin with the eigendecomposition of $\mathbf{K}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$, which does not vary with tuning parameters. After this step, the entire algorithm of NCKQR only involves $\mathcal{O}(n^{2})$ operations. Further details are provided in Section B of the online supplemental material.
When the linear constraints in problem (14) are present, we use the same projection in problem (8) to obtain $\{\tilde{b}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$ as
$\tilde{b}_{\tau_{t}}=b_{\tau_{t}}^{(k)}+\frac{1}{|\hat{S}_{t}|+1}\sum_{i\in\hat{S}_{t}}[y_{i}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}^{(k)}],\ \tilde{\boldsymbol{\alpha}}_{\tau_{t}}=\mathbf{K}^{-1}\boldsymbol{\theta},$ (19)
where $|\hat{S}_{t}|$ is the cardinality of $\hat{S}_{t}$, and $\boldsymbol{\theta}\in\mathbb{R}^{n}$ with $\theta_{i}=y_{i}-\tilde{b}_{\tau_{t}}$ if $i\in\hat{S}$ and $\theta_{i}=\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}^{(k)}$ otherwise. We then use $\{\tilde{b}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$ in place of $\{{b}^{[k]}_{\tau_{t}},{\boldsymbol{\alpha}}^{[k]}_{\tau_{t}}\}_{t=1}^{T}$ in problem (17) to proceed with the algorithm.
We also note that the MM algorithm requires $\gamma\leq\eta$, where $\eta=10^{-5}$. This condition guarantees the proper majorization within the MM algorithm but it may also drive the algorithm overly conservative, slowing progress toward the solution. In our implementation, we begin with $\gamma=\eta=1$ and iteratively reduce both parameters to a quarter of their previous values, terminating the algorithm if the KKT conditions of problem (12) are met. If the algorithm does not terminate when we reach some $\gamma=\eta<10^{-5}$, we maintain $\eta=10^{-5}$ and continue to decrease $\gamma$. Similar to the single-level KQR, the algorithm typically stops after three or four iterations of updating $\gamma$.
The NCKQR algorithm is summarized in Algorithm 2 in the online supplemental material.
4 Numerical Studies
We now use simulation experiments to showcase the performance of our fastkqr algorithm for the single-level KQR and multi-level NCKQR, in Sections 4.1 and 4.2, respectively. Additional simulations and benchmark data applications are present in the online supplemental materials.
4.1 Kernel quantile regression
We compare our fastkqr algorithm with the R package kernlab (Karatzoglou et al., 2004) and the two generic optimizers nlm and optim in the R package stats. We chose the radial basis kernel and explored various combinations of the sample size $n$ and dimension $p$. We considered three different quantile levels: $\tau=0.1$, $0.5$, and $0.9$. For each training data, we applied the four solvers to fit KQR over 50 $\lambda$ values. The optimal tuning parameter $\lambda$ was selected using five-fold cross-validation, and the whole run time was recorded. The selected $\lambda$ was then used to compute the objective value of problem (2). The reported results represent averages from 20 independent repetitions. All computations were carried out on an Intel Core i5 (2.30 GHz) processor.
We use the same simulation model in Friedman et al. (2010). Predictors are generated from $\mathrm{N}(0,1)$, where each pair is correlated with $\rho=0.1$. The response values were generated by
$Y=\sum_{j=1}^{p}X_{j}\beta_{j}+cZ,$ (20)
where $\beta_{j}=(-1)^{j}\exp(-\frac{j-1}{10})$, $Z\sim\mathrm{N}(0,1)$, and $c$ is set so that the signal-to-noise ratio is 3.0.
$\tau$ $n$ fastkqr kernlab nlm optim
0.1 200 obj $0.553_{(0.091)}$ $0.553_{(0.091)}$ $0.555_{(0.087)}$ $0.598_{(0.048)}$
time 3.06 98.01 444.90 885.11
500 obj $0.591_{(0.048)}$ $0.591_{(0.048)}$ $0.592_{(0.048)}$ $0.611_{(0.044)}$
time 24.75 483.48 6669.63 13798.18
1000 obj $0.294_{(0.052)}$ $0.294_{(0.052)}$ $0.294_{(0.051)}$ $0.332_{(0.061)}$
time 114.97 2168.42 42013.01 $>24$h
0.5 200 obj $1.380_{(0.116)}$ $1.380_{(0.116)}$ $1.381_{(0.115)}$ $1.400_{(0.097)}$
time 2.63 106.75 423.24 858.95
500 obj $0.687_{(0.155)}$ $0.687_{(0.155)}$ $0.690_{(0.146)}$ $0.771_{(0.177)}$
time 24.44 593.39 6096.10 12350.00
1000 obj $0.868_{(0.270)}$ $0.868_{(0.270)}$ $0.869_{(0.270)}$ $0.954_{(0.301)}$
time 111.35 2382.14 43142.23 $>24$h
0.9 200 obj $0.522_{(0.119)}$ $0.522_{(0.119)}$ $0.527_{(0.111)}$ $0.577_{(0.076)}$
time 2.78 90.69 556.54 1124.95
500 obj $0.260_{(0.082)}$ $0.260_{(0.082)}$ $0.267_{(0.073)}$ $0.314_{(0.089)}$
time 23.33 445.50 6610.01 13333.11
1000 obj $0.304_{(0.051)}$ $0.304_{(0.051)}$ $0.306_{(0.048)}$ $0.398_{(0.202)}$
time 124.36 2282.20 43501.72 $>24$h
Table 1: Objective values and computation time of four kernel quantile regression solvers for simulation data with $p=5000$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses.
Tables 1 displays the average objective values and computation time for the four solvers, with all the run time including both model training and tuning parameter selection. Our algorithm fastkqr consistently outperforms the other solvers in speed across all the examples. Specifically, with a sample size $n=200$ and dimension $p=5000$, fastkqr outpaces kernlab by a factor of thirty. We also note that fastkqr and kernlab roughly have the same objective values, which are significantly lower than the other two generic optimizers.
4.2 Non-crossing kernel quantile regression
In this section, we compare fastkqr with the R package cvxr (Fu et al., 2020), and nlm and optim for fitting NCKQR. In each scenario, we simultaneously fit three quantile curves with $\tau=0.1,0.5$, and $0.9$ using the four solvers. With 20 independent runs, we calculate the average objective values of problem (12) and the total run time across 50 $\lambda_{2}$ values.
$p$ $n$ fastkqr cvxr nlm optim
100 200 obj $1.307_{(0.132)}$ $1.361_{(0.261)}$ $1.818_{(1.537)}$ $3.497_{(0.568)}$
time 4.04 7138.04 6573.51 13071.21
500 obj $2.529_{(0.127)}$ $2.612_{(0.331)}$
time 31.88 12739.79 $>24$h $>24$h
1000 obj $3.380_{(0.114)}$
time 222.12 $>24$h $>24$h $>24$h
1000 200 obj $1.121_{(0.102)}$ $1.166_{(0.225)}$ $3.219_{(1.414)}$ $4.277_{(0.567)}$
time 2.48 7067.55 6464.77 13036.94
500 obj $2.343_{(0.119)}$ $2.409_{(0.287)}$
time 28.61 15596.45 $>24$h $>24$h
1000 obj $3.372_{(0.132)}$
time 138.58 $>24$h $>24$h $>24$h
5000 200 obj $1.118_{(0.121)}$ $1.163_{(0.257)}$ $3.145_{(1.512)}$ $4.322_{(0.702)}$
time 3.24 7098.27 6328.77 12838.61
500 obj $2.290_{(0.134)}$ $2.352_{(0.278)}$
time 19.56 13180.66 $>24$h $>24$h
1000 obj $3.375_{(0.114)}$
time 117.09 $>24$h $>24$h $>24$h
Table 2: Objective values and computation time of four NCKQR solvers for simulation data with $p=\{100,1000,5000\}$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs, and the standard errors are presented in the parentheses. A result shown as a star “*" means the corresponding solver cannot output a solution due to some numerical issue.
Table 2 presents the average objective values and computation time. Our algorithm is the fastest and the most accurate. For example, when $n=200$, our algorithm took only four seconds, in contrast to cvxr which spent about two hours. Furthermore, when $n$ is increased to 1000, our algorithm spent just 200 seconds, while the other three competitors took over 24 hours.
5 Discussions
In this paper, we have developed fastkqr, a fast algorithm for computing the exact solution of kernel quantile regression. Our approach is based on a finite smoothing algorithm and accelerated proximal gradient descent, further enhanced by a fast spectral technique that optimizes matrix operations. Notably, fastkqr is faster than kernlab while maintaining nearly identical accuracy.
Furthermore, we have addressed the crossing of quantile curves. We have introduced the non-crossing kernel quantile regression with a soft non-crossing penalty, and have expanded fastkqr with a specialized MM algorithm featuring two majorization steps. We show that fastkqr significantly outperforms cvxr in both computational speed and accuracy.
To broaden the applicability of fastkqr to large-scale data analysis, we propose to integrate various kernel approximation techniques into our existing algorithmic framework. Methods such as random features (Rahimi and Recht, 2007) or Nyström subsampling (Rudi et al., 2015) could be employed within the exact update formula of kernel quantile regression to create a cost-effective surrogate of the kernel matrix. These approximation strategies are expected to further enhance the efficiency of our algorithm. We leave a full investigation of this direction for future work.
SUPPLEMENTARY MATERIAL
Appendix A Algorithm
The complete algorithm for solving the KQR in problem (2) is detailed in Algorithm 1. Algorithm 2 summarizes the entire algorithm for computing the NCKQR.
Algorithm 1 The fastkqr algorithm for solving the kernel quantile regression in Problem (2)
1:$\mathbf{y}$, $\mathbf{K}$, $\tau$, and $\lambda^{[1]}>\lambda^{[2]}>\ldots>\lambda^{[L]}$.
2:Carry out the eigen-decomposition $\mathbf{K}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$ only once.
3:Initialize $\gamma=1$, $\delta=1/4$, and $(\tilde{b},\tilde{\boldsymbol{\alpha}})$.
4:for $l=1,2,\ldots,L$ do
5:   Set $\lambda\leftarrow\lambda^{[l]}$.
6:   repeat
7:      Compute $\mathbf{v}=\mathbf{U}\boldsymbol{\Lambda}\boldsymbol{\Pi}_{\gamma,\lambda}^{-1}\mathbf{U}^{\top}\mathbf{1}$ and $g=\bigl{(}n-\mathbf{1}^{\top}\mathbf{U}\boldsymbol{\Lambda}\boldsymbol{\Pi}_{\gamma,\lambda}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\mathbf{1}\bigr{)}^{-1}$, where $\boldsymbol{\Pi}_{\gamma,\lambda}=\boldsymbol{\Lambda}^{2}+2n\gamma\lambda\boldsymbol{\Lambda}.$
8:      Set $\hat{S}\leftarrow\emptyset$.
9:      repeat
10:         Set $k\leftarrow 1$ and $c_{1}\leftarrow 1$.
11:         Initialize $({b}^{[0]},{\boldsymbol{\alpha}}^{[0]})=({b}^{[1]},{\boldsymbol{\alpha}}^{[1]})\leftarrow(\tilde{b},\tilde{\boldsymbol{\alpha}})$.
12:         repeat
13:            Compute $c_{k+1}\leftarrow\frac{1+\sqrt{1+4c_{k}^{2}}}{2}$.
14:            Update
$\left(\begin{array}[]{c}\bar{b}^{(k+1)}\\ \bar{\boldsymbol{\alpha}}^{(k+1)}\end{array}\right)\leftarrow\left(\begin{array}[]{c}b^{(k)}\\ \boldsymbol{\alpha}^{(k)}\end{array}\right)+\left(\frac{c_{k}-1}{c_{k+1}}\right)\left(\begin{array}[]{c}b^{(k)}-b^{(k-1)}\\ \boldsymbol{\alpha}^{(k)}-\boldsymbol{\alpha}^{(k-1)}\end{array}\right).$
15:            Update $\overline{{z}}_{i}\leftarrow\frac{1}{n}{}H_{\gamma,\tau}^{\prime}(y_{i}-\bar{b}^{(k+1)}-\mathbf{K}_{i}^{\top}\boldsymbol{\bar{\alpha}}^{(k+1)}),\ i=1,\ldots,n.$
16:            Calculate from right to left
$\boldsymbol{\mu}\leftarrow g\{\mathbf{1}^{\top}\bar{\mathbf{z}}-\mathbf{v}^{\top}\mathbf{K}(\bar{\mathbf{z}}+n\lambda\bar{\boldsymbol{\alpha}}^{k+1})\}\left(\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\right)+\left(\begin{array}[]{c}0\\ \mathbf{U}\boldsymbol{\Pi}_{\gamma,\lambda}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}(\bar{\mathbf{z}}+n\lambda\bar{\boldsymbol{\alpha}}^{k+1})\end{array}\right).$
17:            Update
$\left(\begin{array}[]{c}b^{(k+1)}\\ \boldsymbol{\alpha}^{(k+1)}\end{array}\right)\leftarrow\left(\begin{array}[]{c}\bar{b}^{(k+1)}\\ \boldsymbol{\bar{\alpha}}^{(k+1)}\end{array}\right)+\boldsymbol{\mu}.$
18:            Update $k\leftarrow{}k+1$.
19:         until the convergence criterion is met.
20:         Update $(\tilde{b},\tilde{\boldsymbol{\alpha}})$ from Problem (8).
21:         Update $\hat{S}\leftarrow\{i:-\gamma\leq y_{i}-\tilde{b}-\mathbf{K}_{i}^{\top}\tilde{\boldsymbol{\alpha}}\leq\gamma\}$.
22:      until the set $\hat{S}$ is unchanged.
23:      Update $\gamma\leftarrow\delta\gamma$.
24:   until the KKT conditions of the KQR problem are satisfied.
25:   Set $(\hat{b}^{[l]},\hat{\boldsymbol{\alpha}}^{[l]})\leftarrow(\tilde{b},\tilde{\boldsymbol{\alpha}})$.
26:KQR solution, $(\hat{b}^{[1]},\hat{\boldsymbol{\alpha}}^{[1]}),(\hat{b}^{[2]},\hat{\boldsymbol{\alpha}}^{[2]}),\ldots,(\hat{b}^{[L]},\hat{\boldsymbol{\alpha}}^{[L]})$.
Algorithm 2 The fastkqr algorithm for solving NCKQR in Problem (12)
1:$\mathbf{y}$, $\mathbf{K}$, $\tau_{1}<\tau_{2}<\ldots<\tau_{T}$, $\lambda_{1}^{[1]}>\lambda_{1}^{[2]}>\ldots>\lambda_{1}^{[L_{1}]}$, and $\lambda_{2}^{[1]}>\lambda_{2}^{[2]}>\ldots>\lambda_{2}^{[L_{2}]}$.
2:Carry out the eigen-decomposition $\mathbf{K}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$ only once.
3:Initialize $\gamma=1$, $\delta=1/4$, and $\{\tilde{b}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$.
4:for $l_{1}=1,2,\ldots,L_{1}$ do
5:   for $l_{2}=1,2,\ldots,L_{2}$ do
6:      Set $\lambda_{1}\leftarrow\lambda_{1}^{[l_{1}]}$ and $\lambda_{2}\leftarrow\lambda_{2}^{[l_{2}]}$.
7:      repeat
8:         Compute $\mathbf{v}=\left(4\lambda_{2}n+1\right)\mathbf{U}\boldsymbol{\Lambda}\boldsymbol{\Pi}_{\gamma,\lambda_{2}}^{-1}\mathbf{U}^{\top}\mathbf{1}$.
9:         Compute $\boldsymbol{\Pi}_{\gamma,\lambda_{2}}=\left(4\lambda_{2}n+1\right)\boldsymbol{\Lambda}\boldsymbol{\Lambda}+\lambda_{2}\varepsilon n+2n\lambda_{2}\gamma\boldsymbol{\Lambda}$.
10:         Compute $g=[\left(4\lambda_{2}n+1\right)n+\lambda_{2}\varepsilon n-\left(4\lambda_{2}n+1\right)^{2}\mathbf{1}^{\top}\mathbf{U}\boldsymbol{\Lambda}\boldsymbol{\Pi}_{\gamma,\lambda_{2}}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\mathbf{1}]^{-1}$.
11:         Set $\hat{S}_{1}=\hat{S}_{2}=\ldots=\hat{S}_{T}\leftarrow\emptyset$.
12:         repeat
13:            Set $k=1$.
14:            Initialize $\{b_{\tau_{t}}^{(k)},\boldsymbol{\alpha}_{\tau_{t}}^{(k)}\}_{t=1}^{T}\leftarrow\{\tilde{b}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$.
15:            repeat
16:               Update $(\mathbf{z}_{\tau_{t}})_{i}\leftarrow H_{\gamma,\tau_{t}}^{\prime}(y_{i}-b^{(k+1)}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k+1)}_{\tau_{t}})$ for each $t=1,2,\ldots,T$.
17:               Update $(\mathbf{q}_{\tau_{t}})_{i}\leftarrow V^{\prime}(\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}+b^{(k)}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}^{(k)}_{\tau_{t+1}}-b^{(k)}_{\tau_{t+1}})$ for each $t=1,2,\ldots,T-1$.
18:               Calculate the update formulae (LABEL:eq:nckqr-formula1), (24) and (LABEL:eq:nckqr-formula3) from right to left.
19:               Update $\{b_{\tau_{t}}^{(k)},\boldsymbol{\alpha}_{\tau_{t}}^{(k)}\}_{t=1}^{T}$ using formula (18).
20:               Update $k\leftarrow{}k+1$.
21:            until the convergence criterion is met.
22:            Update $\{\tilde{b}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$ using formula (19).
23:            Update $\hat{S}_{t}\leftarrow\{i:-\gamma{}\leq{}y_{i}-\tilde{b}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\tilde{\boldsymbol{\alpha}}_{\tau_{t}}\leq\gamma\}$ for each $t=1,2,\ldots,T$.
24:         until all the sets $\hat{S}_{1},\hat{S}_{2},\ldots,\hat{S}_{T}$ are unchanged.
25:         Update $\gamma\leftarrow\delta\gamma$.
26:      until the KKT conditions of the NCKQR problem are satisfied.
27:      Set $\{\hat{b}_{\tau_{t}}^{[l_{1},\ l_{2}]},\hat{\boldsymbol{\alpha}}_{\tau_{t}}^{[l_{1},\ l_{2}]}\}_{t=1}^{T}\leftarrow\{\tilde{b}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}_{\tau_{t}}\}_{t=1}^{T}$.
28:NCKQR solution, $\{\hat{b}_{\tau_{t}}^{[l_{1},\ l_{2}]},\hat{\boldsymbol{\alpha}}_{\tau_{t}}^{[l_{1},\ l_{2}]}\}_{t=1}^{T}$, for each $l_{1}=1,2,\ldots,L_{1}$ and $l_{2}=1,2,\ldots,L_{2}$.
Appendix B The fast spectral technique for NCKQR
In this section, we study the fast spectral technique for the computation of NCKQR in Section 3.3. Specifically, the fast spectral technique is developed to address the repeated computation of inverting $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}\in\mathbb{R}^{(n+1)\times(n+1)}$, in equation (18), for varying values of $\gamma$, $\lambda_{1}$ and $\lambda_{2}$. We calculate $\boldsymbol{\Pi}_{\gamma,\lambda_{1},\lambda_{2}}=\left(4\lambda_{1}n+1\right)\boldsymbol{\Lambda}\boldsymbol{\Lambda}+\lambda_{1}\varepsilon n+2n\lambda_{2}\gamma\boldsymbol{\Lambda}$ for each $\lambda_{1}$ and $\gamma$. By employing $g=1/[(4\lambda_{1}n+1)n+\lambda_{1}\varepsilon n-\left(4\lambda_{1}n+1\right)^{2}\mathbf{1}^{\top}\mathbf{U}\boldsymbol{\Lambda}\boldsymbol{\Pi}_{\gamma,\lambda_{1}}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\mathbf{1}]$ and $\mathbf{v}=\left(4\lambda_{1}n+1\right)\mathbf{U}\boldsymbol{\Lambda}\boldsymbol{\Pi}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\mathbf{U}^{\top}\mathbf{1}$, we can readily attain the desired decomposition
$\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}$ $=\left(\begin{array}[]{cc}n+4\lambda_{1}n^{2}+\lambda_{1}\varepsilon n&\left(4\lambda_{1}n+1\right)\mathbf{1}^{\top}\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}\\ \left(4\lambda_{1}n+1\right)\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}\mathbf{1}&\mathbf{U}\boldsymbol{\Pi}_{\gamma,\lambda_{1},\lambda_{2}}\mathbf{U}^{\top}\end{array}\right)^{-1}$
$=g\left(\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\right)\left(\begin{array}[]{ll}1&-\mathbf{v}^{\top}\end{array}\right)+\left(\begin{array}[]{cc}0&\mathbf{0}^{\top}\\ \mathbf{0}&\mathbf{U}\boldsymbol{\Pi}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\mathbf{U}^{\top}\end{array}\right).$
Rather than computing $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}$, we directly compute the following matrix-vector multiplications.
  • Case 1. When $t=1$, we have
    $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}^{\top}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}\end{array}\right)$ (21)
    $=g\left\{\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}n-\mathbf{v}^{\top}\mathbf{K}\left(\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{2}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t}}\right)\right\}\left(\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\right)$
    $\quad+\left(\begin{array}[]{c}0\\ \mathbf{U}\boldsymbol{\Pi}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\left(\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{2}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t}}\right)\end{array}\right).$
  • Case 2. When $12\leq t\leq T-$ 1, we have
    $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}^{\top}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\end{array}\right)$ (24)
    $=g\left\{\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}-\mathbf{v}^{\top}\mathbf{K}\left(\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{2}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t-1}}\right)\right\}$
    $\quad\cdot\left(\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\right)+\left(\begin{array}[]{c}0\\ \mathbf{U}\boldsymbol{\Pi}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\left(\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{2}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}-n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t-1}}\right)\end{array}\right).$ (29)
  • Case 3. When $t=T$, we have
    $\boldsymbol{\Sigma}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\left(\begin{array}[]{c}\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\\ \mathbf{K}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}-n\lambda_{2}\mathbf{K}^{\top}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{K}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\end{array}\right)$ (30)
    $=g\left\{\mathbf{1}^{\top}\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{1}^{\top}\mathbf{q}^{(k)}_{\tau_{t-1}}\mathbf{v}^{\top}\mathbf{K}\left(\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{2}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t-1}}\right)\right\}\left(\begin{array}[]{c}1\\ -\mathbf{v}\end{array}\right)$
    $\quad+\left(\begin{array}[]{c}0\\ \mathbf{U}\boldsymbol{\Pi}_{\gamma,\lambda_{1},\lambda_{2}}^{-1}\boldsymbol{\Lambda}\mathbf{U}^{\top}\left(\mathbf{z}^{(k)}_{\tau_{t}}+n\lambda_{2}\boldsymbol{\alpha}^{(k)}_{\tau_{t}}+n\lambda_{1}\mathbf{q}^{(k)}_{\tau_{t-1}}\right)\end{array}\right).$
Appendix C Additional Simulations
Table 3 presents the objective values and computation times using simulation data for $p=100$. Notably, fastkqr consistently emerges as the fastest solver, outperforming the other three solvers by at least an order of magnitude, while also achieving the lowest objective value.
$\tau$ $n$ fastkqr kernlab nlm optim
0.1 200 obj $0.551_{(0.077)}$ $0.551_{(0.077)}$ $0.558_{(0.066)}$ $0.583_{(0.051)}$
time 0.85 13.09 376.21 788.85
500 obj $0.553_{(0.045)}$ $0.553_{(0.045)}$ $0.566_{(0.039)}$ $0.588_{(0.037)}$
time 5.74 120.46 5351.71 10447.53
1000 obj $0.558_{(0.025)}$ $0.558_{(0.025)}$ $0.576_{(0.022)}$ $0.591_{(0.024)}$
time 20.91 695.72 45076.91 $>24$h
0.5 200 obj $1.292_{(0.093)}$ $1.292_{(0.093)}$ $1.297_{(0.089)}$ $1.332_{(0.081)}$
time 0.80 12.17 319.15 686.34
500 obj $1.287_{(0.054)}$ $1.287_{(0.054)}$ $1.301_{(0.050)}$ $1.331_{(0.049)}$
time 6.50 153.95 5356.08 10480.23
1000 obj $1.305_{(0.033)}$ $1.305_{(0.033)}$ $1.320_{(0.031)}$ $1.344_{(0.034)}$
time 17.97 686.88 46135.44 $>24$h
0.9 200 obj $0.541_{(0.072)}$ $0.541_{(0.072)}$ $0.547_{(0.065)}$ $0.579_{(0.054)}$
time 0.83 11.67 383.48 798.80
500 obj $0.541_{(0.050)}$ $0.541_{(0.050)}$ $0.557_{(0.040)}$ $0.584_{(0.034)}$
time 5.04 103.57 5156.50 10968.19
1000 obj $0.559_{(0.028)}$ $0.559_{(0.028)}$ $0.576_{(0.023)}$ $0.591_{(0.026)}$
time 22.14 709.62 44067.48 $>24$h
Table 3: Objective values and computation time of four kernel quantile regression solvers for simulation data with $p=100$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses.
Following Yuan (2006), two-dimensional data are generated based on
$Y=\frac{40\exp\left[8\left\{(X_{1}-.5)^{2}+(X_{2}-.5)^{2}\right\}\right]}{\exp\left[8\left\{(X_{1}-.2)^{2}+(X_{2}-.7)^{2}\right\}\right]+\exp\left[8\left\{(X_{1}-.7)^{2}+(X_{2}-.2)^{2}\right\}\right]}+\epsilon,$ (31)
where $X_{1}$ and $X_{2}$ are drawn from Uniform $(0,1)$, and the error term $\epsilon$ is from the standard normal distribution.
Tables 4 presents the average objective values and computation time for the four solvers applied to Example 1, with all the runtime including both model training and tuning parameter selection through five-fold cross-validation. Our algorithm, fastkqr, consistently outperforms the other solvers in speed across all the examples. For instance, when $n=1000$, our fastkqr algorithm was more than an order of magnitude faster than kernlab, and notably, more than 400 times faster than nlm and optim. As the sample size $n$ grows, the computational advantages of our algorithm over the others become more significant. Importantly, despite the significant gains in efficiency, our algorithm continues to deliver objective values that are nearly identical to those of kernlab.
$\tau$ $n$ fastkqr kernlab nlm optim
0.1 200 obj $0.442_{(0.030)}$ $0.442_{(0.030)}$ $0.442_{(0.030)}$ $0.446_{(0.030)}$
time 3.84 11.19 331.47 676.90
500 obj $0.426_{(0.020)}$ $0.426_{(0.020)}$ $0.426_{(0.020)}$ $0.428_{(0.019)}$
time 20.71 102.93 4690.49 9756.53
1000 obj $0.421_{(0.016)}$ $0.421_{(0.016)}$ $0.421_{(0.016)}$ $0.422_{(0.015)}$
time 73.83 814.82 39483.72 $>24$h
0.5 200 obj $1.031_{(0.053)}$ $1.031_{(0.053)}$ $1.032_{(0.053)}$ $1.035_{(0.052)}$
time 3.96 11.40 322.66 701.60
500 obj $0.985_{(0.035)}$ $0.984_{(0.035)}$ $0.985_{(0.035)}$ $0.986_{(0.034)}$
time 23.92 112.55 4814.45 10202.31
1000 obj $0.960_{(0.026)}$ $0.960_{(0.026)}$ $0.960_{(0.026)}$ $0.961_{(0.026)}$
time 76.64 715.08 39854.58 $>24$h
0.9 200 obj $0.468_{(0.042)}$ $0.468_{(0.042)}$ $0.469_{(0.041)}$ $0.475_{(0.039)}$
time 3.80 9.79 298.61 704.44
500 obj $0.457_{(0.020)}$ $0.457_{(0.020)}$ $0.457_{(0.020)}$ $0.459_{(0.020)}$
time 22.93 110.13 4116.96 9816.24
1000 obj $0.445_{(0.017)}$ $0.445_{(0.017)}$ $0.446_{(0.017)}$ $0.447_{(0.017)}$
time 76.01 778.49 38489.13 $>24$h
Table 4: Objective values and computation time of four kernel quantile regression solvers using Yuan (2006) with $p=2$, $n=\{200,500,1000\}$, and $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs, and the standard errors are presented in the parentheses.
Appendix D Benchmark Data Applications
We used four benchmark data from the R packages MASS and mlbench to further compare our fastkqr with the three competitors: kernlab, nlm, and optim. The first data set crabs includes five morphometric measurements from 50 crabs of the Leptograpsus variegatus species, collected in Fremantle, Western Australia. For our analysis, categorical variables were converted to dummy variables, with carapace width as the response variable, and variable “index” was removed. The second data set GAGurine (GAG) was introduced in Section 1. The third data set, BostonHousing (BH), provides housing information in the Boston area with the median value of owner-occupied homes as the predictor variable. The last data set, geyser, records details of the “Old Faithful" geyser in the Yellowstone National Park, using eruption time as the predictor variable.
For each benchmark data, we employed the five-fold cross-validation to select the optimal tuning parameter, which was then used to calculate the objective values for accuracy assessment. We also recorded the total computation time, including both the model training and parameter tuning via the five-fold cross-validation. Notably, fastkqr consistently demonstrated superior efficiency compared to kernlab. Both fastkqr and kernlab achieved the highest accuracy. The computational advantages of fastkqr become increasingly evident with larger sample sizes. For instance, when fitting the mcycle data with 133 observations, fastkqr required only half the time needed by kernlab. For the larger BostonHousing data consisting of 506 observations, fastkqr was ten times faster than cvxr.
$\text{data}_{(n,p)}$ $\tau$ fastkqr kernlab nlm optim
$\text{crabs}_{(200,8)}$ 0.1 obj $0.486_{(0.069)}$ $0.486_{(0.069)}$ $0.487_{(0.069)}$ $1.800_{(0.420)}$
time 3.12 9.38 263.23 731.41
0.5 obj $1.595_{(0.176)}$ $1.595_{(0.176)}$ $1.596_{(0.175)}$ $3.210_{(1.130)}$
time 2.83 10.10 241.36 801.27
0.9 obj $0.476_{(0.051)}$ $0.476_{(0.051)}$ $0.477_{(0.051)}$ $0.994_{(0.899)}$
time 3.10 9.02 261.37 803.47
$\text{GAG}_{(314,1)}$ 0.1 obj $0.685_{(0.026)}$ $0.685_{(0.026)}$ $0.685_{(0.026)}$ $0.737_{(0.044)}$
time 11.23 55.41 1326.94 3190.88
0.5 obj $2.371_{(0.140)}$ $2.371_{(0.140)}$ $2.371_{(0.140)}$ $2.372_{(0.140)}$
time 12.62 46.02 1011.65 2234.01
0.9 obj $1.652_{(0.113)}$ $1.652_{(0.113)}$ $1.652_{(0.113)}$ $1.654_{(0.112)}$
time 11.51 49.63 1080.17 3134.22
$\text{mcycle}_{(133,1)}$ 0.1 obj $4.843_{(0.212)}$ $4.843_{(0.212)}$ $4.851_{(0.212)}$ $5.580_{(0.307)}$
time 3.90 7.55 97.60 236.85
0.5 obj $9.969_{(0.489)}$ $9.969_{(0.489)}$ $9.989_{(0.487)}$ $10.340_{(0.450)}$
time 3.28 4.45 197.67 420.41
0.9 obj $4.210_{(0.527)}$ $4.210_{(0.527)}$ $4.215_{(0.524)}$ $4.469_{(0.444)}$
time 2.66 6.75 125.48 298.21
$\text{BH}_{(506,14)}$ 0.1 obj $0.826_{(0.055)}$ $0.826_{(0.055)}$ $0.827_{(0.055)}$ $1.020_{(0.221)}$
time 16.99 161.32 7473.12 12340.06
0.5 obj $1.941_{(0.056)}$ $1.941_{(0.056)}$ $1.946_{(0.055)}$ $2.079_{(0.118)}$
time 14.13 116.88 5396.74 11370.48
0.9 obj $1.051_{(0.028)}$ $1.051_{(0.028)}$ $1.072_{(0.025)}$ $1.553_{(0.455)}$
time 16.05 152.68 5317.58 10527.89
Table 5: Objective values and computation time of four KQR solvers for benchmark data with $\tau=\{0.1,0.5,0.9\}$. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses.
For the NCKQR problem, we employed the same four benchmark data to compare fastkqr algorithm with its three competitors: cvxr, nlm, and optim. For each benchmark data, we presented the average objective values and the total computation time across a range of $\lambda_{2}$ values and five quantile levels: 0.1, 0.3, 0.5, 0.7, and 0.9. Table 6 show that our algorithm consistently delivers the best accuracy and the computation is the fastest. In all the examples, fastkqr consistently outperforms the other competitors, being at least 80 times faster.
$\text{data}_{(n,p)}$ fastkqr kernlab nlm optim
$\text{crabs}_{(200,8)}$ obj $0.921_{(0.023)}$ $1.051_{(0.593)}$ $7.070_{(10.791)}$ $12.280_{(6.382)}$
time 77.98 6125.56 6590.67 13067.35
$\text{GAG}_{(314,1)}$ obj $5.067_{(0.083)}$ $5.407_{(1.052)}$ $6.312_{(2.497)}$ $6.036_{(1.012)}$
time 31.03 9316.60 18614.09 39213.16
$\text{mcycle}_{(133,1)}$ obj $30.575_{(0.397)}$ $32.025_{(5.123)}$ $35.535_{(9.088)}$ $38.662_{(5.332)}$
time 10.40 6736.39 2683.09 5990.27
$\text{BH}_{(506,14)}$ obj $4.006_{(0.044)}$ $5.423_{(4.243)}$
time 321.08 17850.87 $>24$h $>24$h
Table 6: Objective values and computation time of four NCKQR solvers for benchmark data. The numbers are the average quantities over 20 independent runs and the standard errors are presented in the parentheses. A result shown as a star “*" means the corresponding solver cannot output a solution due to some numerical issue.
Appendix E Technical Proofs
In this section, we provide all technical proofs for the results presented in the main article.
E.1 Proof of Proposition 1
To establish the proposition, first we prove the following lemma:
Lemma 8.
Let $(b^{\gamma},\boldsymbol{\alpha}^{\gamma})=\operatorname*{argmin}_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}G^{\gamma}(b,\boldsymbol{\alpha})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}$, then we have
$G(b^{\gamma},\boldsymbol{\alpha}^{\gamma})-\frac{\gamma}{4}\leq{}G(\hat{b},\hat{\boldsymbol{\alpha}})\leq{}G(b^{\gamma},\boldsymbol{\alpha}^{\gamma}).$
Proof.
Recall that
$H_{\gamma,\tau}(t)=\begin{cases}(\tau-1)t&\text{if}\enskip{}t<-\gamma,\\ \frac{t^{2}}{4\gamma}+t(\tau-\frac{1}{2})+\frac{\gamma}{4}&\text{if}\enskip{}-\gamma\leq t\leq\gamma,\\ \tau{}t&\text{if}\enskip{}t>\gamma,\end{cases}$
and
$\rho_{\tau}(t)=\begin{cases}(\tau-1)t&\text{if}\enskip{}t\leq 0,\\ \tau{}t&\text{if}\enskip{}t>0.\end{cases}$
Consider the following three cases:
  • Case 1. When $t<-\gamma<0$ or $t>\gamma>0$, we have
    $H_{\tau,\gamma}(t)-\rho_{\tau}(t)=0.$
  • Case 2. When $-\gamma\leq t\leq 0$, we have
    $0{}\leq{}H_{\tau,\gamma}(t)-\rho_{\tau}(t)=\frac{1}{4\gamma}(t+\gamma)^{2}{}\leq{}\frac{\gamma}{4}.$
  • Case 3. When $0<t\leq\gamma$, we have
    $0{}\leq{}H_{\tau,\gamma}(t)-\rho_{\tau}(t)=\frac{1}{4\gamma}(t-\gamma)^{2}{}<{}\frac{\gamma}{4}.$
Thus, for any $t\in\mathbb{R}$, $40\leq{}H_{\tau,\gamma}(t)-\rho_{\tau}(t)\leq\gamma/$ 4. It follows that
$0\leq{}G^{\gamma}(b,\boldsymbol{\alpha})-G(b,\boldsymbol{\alpha})\leq\frac{\gamma}{4},\forall{}b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n},$ (32)
which implies $G(b^{\gamma},\boldsymbol{\alpha}^{\gamma})\leq{}G^{\gamma}(b^{\gamma},\boldsymbol{\alpha}^{\gamma})$ and $G^{\gamma}(\hat{b},\hat{\boldsymbol{\alpha}})\leq{}G(\hat{b},\hat{\boldsymbol{\alpha}})+\gamma/4$. By optimality of $(\hat{b},\hat{\boldsymbol{\alpha}})$ and $(b^{\gamma},\boldsymbol{\alpha}^{\gamma})$, we have
$G(\hat{b},\hat{\boldsymbol{\alpha}})\leq{}G(b^{\gamma},\boldsymbol{\alpha}^{\gamma})\leq{}G^{\gamma}(b^{\gamma},\boldsymbol{\alpha}^{\gamma})\leq{}G^{\gamma}(\hat{b},\hat{\boldsymbol{\alpha}})\leq{}G(\hat{b},\hat{\boldsymbol{\alpha}})+\frac{\gamma}{4}.$
Now consider the following constrained problem
$\begin{split}&\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}G^{\gamma}(b,\boldsymbol{\alpha})\\ &\text{subject to}\enskip{}y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in S_{0}.\end{split}$ (33)
The Lagrangian of problem (33) is
$L(b,\boldsymbol{\alpha},\xi_{i})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,{}\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}+\sum_{i\in S_{0}}\xi_{i}(b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}-y_{i}),$ (34)
where the $\xi_{i}$’s are the Lagrangian multipliers. By the optimality of $(\hat{\boldsymbol{\alpha}}^{\gamma},\hat{b}^{\gamma})$ in problem (33), we have
$\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in S_{0}}\xi_{i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in S_{0}}\xi_{i}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\,i\in S_{0},\end{array}\right.$ (35)
where $H^{\prime}_{\gamma,\tau}(\cdot)$ denotes the derivative of $H_{\gamma,\tau}(\cdot)$. Note that the subdifferential of $\rho_{\tau}(\cdot)$ is
$\partial\rho_{\tau}(t)=\begin{cases}\{\tau-1\},&\text{if}\enskip{}t<0,\\ [\tau-1,\tau],&\text{if}\enskip{}t=0,\\ \{\tau\},&\text{if}\enskip{}t>0.\end{cases}$
According to the definition of $H_{\gamma,{}\tau}(t)$, one has $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\notin S_{0}$ and $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\tau-1/2{}\in{}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\in S_{0}$. Therefore, it follows from expression (35) that
$\left\{\begin{array}[]{l}\mathbf{0}\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in S_{0}}\xi_{i}\mathbf{K}_{i},\\ 0\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in S_{0}}\xi_{i}.\end{array}\right.$ (36)
We can see that $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ satisfies the KKT conditions of the convex problem:
$\begin{split}&\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}\\ &\text{subject to}\enskip{}y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in S_{0}.\end{split}$ (37)
Therefore, $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ is the optimal solution of problem (37). This implies that
$G(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})\leq{}G(\hat{b},\hat{\boldsymbol{\alpha}})\leq{}G(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma}),$
The first inequality is justified because $(\hat{b},\hat{\boldsymbol{\alpha}})$ is a feasible point of problem (37). The second inequality follows since $(\hat{b},\hat{\boldsymbol{\alpha}})$ is the unconstrained minimizer of $G(b,\boldsymbol{\alpha})$. Thus, we have $G(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=G(\hat{b},\hat{\boldsymbol{\alpha}})$. Given the uniqueness of $(\hat{b},\hat{\boldsymbol{\alpha}})$, it follows that $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=(\hat{b},\hat{\boldsymbol{\alpha}})$.
E.2 Proof of Theorem 2
Since $S\subseteq{}S_{0}$, we have $y_{i}=\hat{b}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}$ for $i\in{}S$. Therefore, $(\hat{b},\hat{\boldsymbol{\alpha}})$ is a feasible point of the minimization problem
$\begin{split}\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}&G^{\gamma}(b,\boldsymbol{\alpha})\\ \text{subject to}\enskip{}&y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},i\in{}S.\end{split}$
By the optimality of $(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})$, we have $G^{\gamma}(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})\leq{}G^{\gamma}(\hat{b},\hat{\boldsymbol{\alpha}})$. It then follows from inequality (32) that
$\begin{split}&G(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})-G(\hat{b},\hat{\boldsymbol{\alpha}})\\ &=[G(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})-G^{\gamma}(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})]+[G^{\gamma}(\tilde{b}^{\gamma},\tilde{\boldsymbol{\alpha}}^{\gamma})-G^{\gamma}(\hat{b},\hat{\boldsymbol{\alpha}})]+[G^{\gamma}(\hat{b},\hat{\boldsymbol{\alpha}})-G(\hat{b},\hat{\boldsymbol{\alpha}})]\\ &\leq 0+0+\frac{\gamma}{4}<\frac{\gamma^{*}}{4}\leq\rho.\end{split}$
By the definition of $D_{\gamma_{0}/2}$, we know $(\tilde{\boldsymbol{\alpha}}^{\gamma},\tilde{b}^{\gamma})\notin{}D_{\gamma_{0}/2}$, and therefore,
$|\tilde{b}^{\gamma}+\mathbf{K}_{i}^{\top}\tilde{\boldsymbol{\alpha}}^{\gamma}-\hat{b}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}|<\frac{\gamma_{0}}{2},\forall{}i=1,\ldots,n.$
This implies that for any $i\in\mathcal{E}(S)$,
$|y_{i}-\hat{b}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}|\leq{}|y_{i}-\tilde{b}^{\gamma}-\mathbf{K}_{i}^{\top}\tilde{\boldsymbol{\alpha}}^{\gamma}|+|\tilde{b}^{\gamma}+\mathbf{K}_{i}^{\top}\tilde{\boldsymbol{\alpha}}^{\gamma}-\hat{b}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}|<\gamma+\frac{\gamma_{0}}{2}<\gamma^{*}+\frac{\gamma_{0}}{2}<\gamma_{0},$
which, by the definition of $\gamma_{0}$, further implies that $i\in{}S_{0}$. We conclude that $S\subseteq\mathcal{E}(S)\subseteq{}S_{0}$.
E.3 Proof of Theorem 3
Note that since $S\subseteq{}S_{0}$ and $\gamma\in(0,\gamma^{*})$, by Theorem 2, we have $S\subseteq\tilde{S}_{0}^{\gamma}\subseteq{}S_{0}$. Similarly, for any $j\in\mathbb{N}^{+}$, if $\tilde{S}_{j-1}^{\gamma}\subseteq{}S_{0}$, then by Theorem 2 again, we have $\tilde{S}_{j-1}^{\gamma}\subseteq\tilde{S}_{j}^{\gamma}\subseteq{}S_{0}$. Therefore, by induction, we have
$S\subseteq\tilde{S}_{0}^{\gamma}\subseteq\tilde{S}_{1}^{\gamma}\subseteq\tilde{S}_{2}^{\gamma}\subseteq\cdots\subseteq{}S_{0}.$
Since $S_{0}$ is a finite set, there exists a finite $j^{*}\in\mathbb{N}^{+}$ such that $\tilde{S}_{j^{*}-1}^{\gamma}=\tilde{S}_{j^{*}}^{\gamma}{}\subseteq{}S_{0}.$ At this moment $\tilde{S}_{j^{*}}^{\gamma}$ is the set, $\hat{S}$, we want to find.
Consider the following constrained problem
$\begin{split}\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}&G^{\gamma}(b,\boldsymbol{\alpha})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}\\ \text{subject to}\enskip{}&y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha},\,i\in\hat{S}.\end{split}$ (38)
Note that the minimizer of problem (38) is $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$. Therefore, $y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}$ for $i\in\hat{S}$. The Lagrangian of problem (38) is
$L(b,\boldsymbol{\alpha},\xi_{i})=\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,{}\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}+\sum_{i\in\hat{S}}\xi_{i}(b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}-y_{i}),$ (39)
where the $\xi_{i}$’s are the Lagrangian multipliers. Since $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ is the optimal solution of problem (38), we have
$\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in\hat{S}}\xi_{i}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},i\in\hat{S},\end{array}\right.$ (40)
where $H^{\prime}_{\gamma,\tau}(\cdot)$ denotes the derivative of $H_{\gamma,\tau}(\cdot)$. Recall that $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\notin{\hat{S}}$ and $H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})=\tau-1/2{}\in{}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})$ when $i\in\hat{S}$. Therefore, it follows from the first line of expression (40) that
$\begin{split}\mathbf{0}&=-\frac{1}{n}\sum_{i\notin\hat{S}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}\\ &\in-\frac{1}{n}\sum_{i\notin\hat{S}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}\\ &=-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})\mathbf{K}_{i}+\lambda\mathbf{K}\hat{\alpha}^{\gamma}+\sum_{i\in\hat{S}}\xi_{i}\mathbf{K}_{i}.\end{split}$ (41)
Similarly, it can be shown from the second line of display (40) that
$0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in\hat{S}}\xi_{i}\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma})+\sum_{i\in\hat{S}}\xi_{i}.$ (42)
Now consider the constrained problem
$\begin{split}\min_{b\in\mathbb{R},\boldsymbol{\alpha}\in\mathbb{R}^{n}}\enskip{}&\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha})+\frac{\lambda}{2}\boldsymbol{\alpha}^{\top}\mathbf{K}\boldsymbol{\alpha}\\ \text{subject to}\enskip{}&y_{i}=b+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}+b,\,i\in\hat{S}.\end{split}$ (43)
We can see from expressions (41), (42) and the third line of expression (40) that $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ satisfies the KKT conditions of the convex problem (43). Therefore, $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})$ is the minimizer of problem (43). Moreover, it can be easily seen that $(\hat{b},\hat{\boldsymbol{\alpha}})$ is a feasible point of problem (43) since $\hat{S}\subseteq{}S_{0}$. This implies that
$G(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})\leq{}G(\hat{b},\hat{\boldsymbol{\alpha}})\leq{}G(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma}),$
where the second inequality follows from the fact that $(\hat{b},\hat{\boldsymbol{\alpha}})$ is the unconstrained minimizer of $G(b,\boldsymbol{\alpha})$. Thus, we have $G(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=G(\hat{b},\hat{\boldsymbol{\alpha}})$. By the uniqueness of $(\hat{b},\hat{\boldsymbol{\alpha}})$, we obtain $(\hat{b}^{\gamma},\hat{\boldsymbol{\alpha}}^{\gamma})=(\hat{b},\hat{\boldsymbol{\alpha}})$.
E.4 Proof of Proposition 5
To establish the proposition, first we prove the following lemma:
Lemma 9.
Let $\{b^{\gamma}_{\tau_{t}},\boldsymbol{\alpha}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}=\operatorname*{argmin}_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\\ }Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)$, then we have
$Q\left(\{b^{\gamma}_{\tau_{t}},\boldsymbol{\alpha}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)-\frac{T}{4}\gamma\leq{}Q\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}\right)\leq{}Q\left(\{b^{\gamma}_{\tau_{t}},\boldsymbol{\alpha}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right).$
Proof.
Based on the proof of Proposition 1, we have
$0\leq Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)-Q\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)=\sum^{T}_{t=1}\left[G^{\gamma}(\boldsymbol{\alpha}_{\tau_{t}},b_{\tau_{t}})-G(\boldsymbol{\alpha}_{\tau_{t}},b_{\tau_{t}})\right]\leq\frac{T}{4}\gamma,$ (44)
which implies
$0\leq Q^{\gamma}\left(\{b^{\gamma}_{\tau_{t}},\boldsymbol{\alpha}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)-Q\left(\{b^{\gamma}_{\tau_{t}},\boldsymbol{\alpha}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)\leq\frac{T}{4}\gamma$
and
$Q\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}\right)\leq Q^{\gamma}\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}\right).$
By the optimality of $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}$ and $\{b^{\gamma}_{\tau_{t}},\boldsymbol{\alpha}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$, we have
$\begin{split}&Q\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}\right)\leq{}Q\left(\{b^{\gamma}_{\tau_{t}},\boldsymbol{\alpha}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)\leq{}Q^{\gamma}\left(\{b^{\gamma}_{\tau_{t}},\boldsymbol{\alpha}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)\\ &\leq{}Q^{\gamma}\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}\right)\leq{}Q\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}\right)+\frac{T}{4}\gamma.\end{split}$
Consider the following constrained problem
$\begin{split}&\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)\\ &\text{subject to}\enskip{}y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in S_{0,t},1\leq t\leq T.\end{split}$ (45)
The Lagrangian of problem (45) is
$\begin{split}L(b,\boldsymbol{\alpha},\xi)=&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}^{\top}_{\tau_{t}}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\sum^{T}_{t=1}\sum_{i\in S_{0,t}}\xi_{i,t}(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-y_{i})\\ &+\lambda_{1}\sum^{T-1}_{t=1}\sum_{i=1}^{n}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}}),\end{split}$ (46)
where the $\xi_{i,t}$’s are the Lagrangian multipliers. Let $H^{\prime}_{\gamma,\tau}(\cdot)$ denote the derivative of $H_{\gamma,\tau}(\cdot)$ and $V^{\prime}(\cdot)$ denote the derivative of $V(\cdot)$. Consider the following three cases, by the optimality of $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ in problem (51),
  • Case 1. When $t=1$, we have
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in S_{0,t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{2}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in S_{0,t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\enskip{}i\in S_{0,t}.\end{array}\right.$ (47)
  • Case 2. When $1<t<T$, we have
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in S_{0,t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in S_{0,t}}\xi_{i,t}=0\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\,i\in S_{0,t}.\end{array}\right.$ (48)
  • When $t=T$, we have
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in S_{0,t}}\xi_{t,i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in S_{0,t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma},\,i\in S_{0,t}.\end{array}\right.$ (49)
Therefore, it follows from expression (47) that
$\left\{\begin{array}[]{l}\mathbf{0}\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in S_{0,t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i},\\ 0\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in S_{0,t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\,i\in S_{0,t},\end{array}\right.$
Similarly, it can be shown for the rest two displays (48) and (49) that
$\left\{\begin{array}[]{l}\mathbf{0}\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in S_{0,t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i},\\ 0\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in S_{0,t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},i\in S_{0,t},\end{array}\right.$
and
$\left\{\begin{array}[]{l}\mathbf{0}\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in S_{0,t}}\xi_{i,t}\mathbf{K}_{i}\\ 0\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in S_{0,t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},i\in S_{0,t}.\end{array}\right.$
We can see that $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ satisfies the KKT conditions of the following convex problem:
$\begin{split}\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}_{\tau_{t}}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\lambda_{1}\sum^{T-1}_{t=1}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}})\\ \text{subject to}\enskip{}&y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in S_{0,t},t=1,\ldots,T.\end{split}$ (50)
Therefore, $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ is the minimizer of problem (50). Moreover, since $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}$ is a feasible point of (50), we have
$Q\left(\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)\leq{}Q\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}\right)\leq{}Q\left(\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right),$
where the second inequality follows from the fact that ($(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1})$ is the unconstrained minimizer of $Q\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)$. Thus, we have $Q\left(\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)=Q\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}\right)$ and by the uniqueness of $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}$, we obtain
$\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}=\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1},$
which completes the proof.
E.5 Proof of Theorem 6
Since $S_{t}\subseteq{}S_{0,t}$, we have $y_{i}=\hat{b}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}_{\tau_{t}}$ for $i\in{}S_{t}$ and $1\leq t\leq T$. Therefore, $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}$ is a feasible point of the minimization problem
$\begin{split}\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}&Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)\\ \text{subject to}\enskip{}&y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},i\in{}S_{t},1\leq t\leq T.\end{split}$
By the optimality of $\{\tilde{b}^{\gamma}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$, we have
$Q^{\gamma}\left(\{\tilde{b}^{\gamma}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)\leq{}Q^{\gamma}\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}\right).$
It then follows from (44) that
$\begin{split}&Q\left(\{\tilde{b}^{\gamma}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)-Q\left(\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)\\ &=\left[Q\left(\{\tilde{b}^{\gamma}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)-Q^{\gamma}\left(\{\tilde{b}^{\gamma}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)\right]+\left[Q^{\gamma}\left(\{\tilde{b}^{\gamma}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)-Q^{\gamma}\left(\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)\right]\\ &\quad+\left[Q^{\gamma}\left(\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)-Q\left(\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)\right]\\ &\leq 0+0+\frac{T}{4}\gamma<\frac{T}{4}\gamma^{*}\leq\rho.\end{split}$
By the definition of $D_{\gamma_{0,t}/2}$, we know $(\tilde{b}^{\gamma}_{\tau_{t}},\tilde{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\notin{}D_{\gamma_{0,t}/2}$ for all quantile levels, and therefore,
$|\tilde{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\tilde{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}_{\tau_{t}}|<\frac{\gamma_{0,t}}{2},\forall{}i=1,\ldots,n,t=1,\ldots,T.$
This implies that for any $i\in\mathcal{E}_{t}(S_{1},S_{2},\cdots,S_{T})$ and $1\leq t\leq T$,
$\begin{split}&|y_{i}-\hat{b}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}_{\tau_{t}}|\leq{}|y_{i}-\tilde{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\tilde{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}|+|\tilde{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\tilde{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}_{\tau_{t}}|\\ &<\gamma+\frac{\gamma_{0,t}}{2}<\gamma^{*}+\frac{\gamma_{0,t}}{2}\leq\gamma_{0,t},\end{split}$
which, by the definition of $\gamma_{0,t}$, further implies that $i\in{}S_{0,t}$ Thus, $S_{t}\subseteq\mathcal{E}_{t}(S_{1},S_{2},\cdots,S_{T})\subseteq{}S_{0,t}$.
E.6 Proof of Theorem 7
Note that since $S_{t}\subseteq{}S_{0,t}$ and $\gamma\in(0,\gamma^{*})$, by Theorem 6, we have $S_{t}\subseteq\tilde{S}_{0,t}^{\gamma}\subseteq{}S_{0,t}$ for all $t$. Similarly, for any $j\in\mathbb{N}^{+}$, if $\tilde{S}_{j-1,t}^{\gamma}\subseteq{}S_{0,t}$, then by Theorem 6 again, we have $\tilde{S}_{j-1,t}^{\gamma}\subseteq\tilde{S}_{j,t}^{\gamma}\subseteq{}S_{0,t}$ for all $t$. Therefore, by mathematical induction, we have
$S_{t}\subseteq\tilde{S}_{0,t}^{\gamma}\subseteq\tilde{S}_{1,t}^{\gamma}\subseteq\tilde{S}_{2,t}^{\gamma}\subseteq\cdots\subseteq{}S_{0,t},\quad t=1,\ldots,T.$
Since $S_{0,t}$ is a finite set, there exists a finite $j^{*}\in\mathbb{N}^{+}$ such that $\tilde{S}_{j^{*}-1,t}^{\gamma}=\tilde{S}_{j^{*},t}^{\gamma}{}\subseteq{}S_{0,t}.$ At this moment $\tilde{S}_{j^{*},t}^{\gamma}$ is the target set $\hat{S}_{t}$.
Consider the following constrained problem
$\begin{split}&\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}Q^{\gamma}\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)\\ &\text{subject to}\enskip{}y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in\hat{S}_{t},1\leq t\leq T.\end{split}$ (51)
The Lagrangian of problem (51) is
$\begin{split}L(b,\boldsymbol{\alpha},\xi)=&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}H_{\gamma,\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}^{\top}_{\tau_{t}}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\sum^{T}_{t=1}\sum_{i\in\hat{S}_{t}^{\gamma}}\xi_{i,t}(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-y_{i})\\ &+\lambda_{1}\sum^{T-1}_{t=1}\sum_{i=1}^{n}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}}),\end{split}$ (52)
where the $\xi_{i,t}$’s are the Lagrangian multipliers. Consider the following three cases, by the optimality of $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ in problem (51),
  • Case 1. When $t=1$, we have
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{2}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\enskip{}i\in\hat{S}_{t}.\end{array}\right.$ (53)
  • Case 2. When $1<t<T$, we have
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}=0\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\,i\in\hat{S}_{t}.\end{array}\right.$ (54)
  • When $t=T$, we have
    $\left\{\begin{array}[]{l}-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in\hat{S}_{t}}\xi_{l,i}\mathbf{K}_{i}=\mathbf{0},\\ -\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}=0,\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\,i\in\hat{S}_{t}.\end{array}\right.$ (55)
Therefore, it follows from expression (53) that
$\left\{\begin{array}[]{l}\mathbf{0}=-\frac{1}{n}\sum_{i\notin\hat{S}_{t}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}_{t}}H^{\prime}_{\gamma,\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}\\ \quad\in-\frac{1}{n}\sum_{i\notin\hat{S}_{t}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}-\frac{1}{n}\sum_{i\in\hat{S}_{t}}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}\\ \quad=-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i},\\ 0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},\,i\in\hat{S}_{t},\end{array}\right.$ (56)
Similarly, it can be shown for expressions (54) and (55), we have
$\left\{\begin{array}[]{l}\mathbf{0}=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})\mathbf{K}_{i},\\ 0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad+\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}-\hat{b}^{\gamma}_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t+1}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},i\in\hat{S}_{t},\end{array}\right.$ (57)
and
$\left\{\begin{array}[]{l}\mathbf{0}=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\lambda_{2}\mathbf{K}\hat{\alpha}^{\gamma}_{\tau_{t}}\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\mathbf{K}_{i}+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\mathbf{K}_{i}\\ 0=-\frac{1}{n}\sum_{i=1}^{n}H^{\prime}_{\gamma,{}\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t}\\ \quad\in-\frac{1}{n}\sum_{i=1}^{n}\partial\rho_{\tau}(y_{i}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})\\ \quad\quad-\lambda_{1}\sum_{i=1}^{n}V^{\prime}(\hat{b}^{\gamma}_{\tau_{t-1}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t-1}}-\hat{b}^{\gamma}_{\tau_{t}}-\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}})+\sum_{i\in\hat{S}_{t}}\xi_{i,t},\\ y_{i}=\hat{b}^{\gamma}_{\tau_{t}}+\mathbf{K}_{i}^{\top}\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}},i\in\hat{S}_{t}.\end{array}\right.$ (58)
Now consider the constrained problem
$\begin{split}\min_{\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}}\enskip{}&\sum^{T}_{t=1}\left[\frac{1}{n}\sum_{i=1}^{n}\rho_{\tau}(y_{i}-b_{\tau_{t}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}})+\frac{\lambda_{2}}{2}\boldsymbol{\alpha}_{\tau_{t}}^{\top}\mathbf{K}\boldsymbol{\alpha}_{\tau_{t}}\right]\\ &+\lambda_{1}\sum^{T-1}_{t=1}V(b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}}-b_{\tau_{t+1}}-\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t+1}})\\ \text{subject to}\enskip{}&y_{i}=b_{\tau_{t}}+\mathbf{K}_{i}^{\top}\boldsymbol{\alpha}_{\tau_{t}},\,i\in\hat{S}_{t},t=1,\ldots,T.\end{split}$ (59)
We see that $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ satisfies the KKT conditions of the convex problem (59). Therefore, $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}$ is the minimizer of problem (59). Moreover, it can be easily seen that $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}$ is a feasible point of (59) since $\hat{S}_{t}\subseteq{}S_{0}$. This implies that
$Q\left(\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)\leq{}Q\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}\right)\leq{}Q\left(\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right),$
where the second inequality follows from the fact that ($(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1})$ is the unconstrained minimizer of $Q\left(\{b_{\tau_{t}},\boldsymbol{\alpha}_{\tau_{t}}\}^{T}_{t=1}\right)$. Thus, we have $Q\left(\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}\right)=Q\left(\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}\right)$ and by the uniqueness of $\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}$, we conclude that $\{\hat{b}^{\gamma}_{\tau_{t}},\hat{\boldsymbol{\alpha}}^{\gamma}_{\tau_{t}}\}^{T}_{t=1}=\{\hat{b}_{\tau_{t}},\hat{\boldsymbol{\alpha}}_{\tau_{t}}\}^{T}_{t=1}$.